### v2

# Efficient Shapley Value-based Non-Uniform Pruning of Large Language Models

Abstract
Pruning large language models (LLMs) is a promising solution for reducing model sizes and computational complexity while preserving performance. Traditional layer-wise pruning methods often adopt a uniform sparsity approach across all layers, which leads to suboptimal performance due to the varying significance of individual transformer layers within the model not being accounted for. To this end, we propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs. This approach quantifies the contribution of each transformer layer to the overall model performance, enabling the assignment of tailored pruning budgets to different layers to retain critical parameters. To further improve efficiency, we design the Sliding Window-based Shapley Value approximation method. It substantially reduces computational overhead compared to exact SV calculation methods. Extensive experiments on various LLMs including LLaMA-v1, LLaMA-v2 and OPT demonstrate the effectiveness of the proposed approach. The results reveal that non-uniform pruning significantly enhances the performance of pruned models. Notably, SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity.
1 Introduction
Large language models (LLMs) have emerged as a transformative technology, demonstrating remarkable capabilities in natural language understanding and generation tasks Brown et al. (2020). LLMs have showcased significant adaptability through fine-tuning, enabling their deployment in highly specialized applications. These advantages underscore the critical role of LLMs in solving real-world challenges Patterson et al. (2021). Despite these capabilities, the deployment of LLMs has been hindered by their immense computational demands. Modern LLMs often consist of billions or even trillions of parameters Luccioni et al. (2023), as seen in models like GPT-3 (175 billion parameters) and PaLM (540 billion parameters). The immense scale incurs significant memory, storage and power costs, making it challenging to run these models on resource-constrained devices. To address these issues, researchers have increasingly turned to model compression techniques, particularly pruning, to reduce the size and computational requirements of LLMs, while retaining their performance Frantar and Alistarh (2023).
Pruning LeCun et al. (1989); Han et al. (2015b, a) is one of the most prominent model compression techniques. It aims to remove redundant or less important parameters from neural networks, thereby reducing their sizes and computational complexity. Recent advances in pruning have progressed from unstructured sparsity (i.e., individual weight removal) to structured sparsity (i.e., eliminating entire neurons, heads or layers), thereby enabling hardware-efficient implementation Frantar and Alistarh (2023); Jaiswal et al. (2024a); Ma et al. (2023); Xia et al. (2023). For LLMs, several state-of-the-art methods have been proposed, including magnitude pruning, lottery ticket hypothesis approaches, and structured pruning based on attention mechanisms. These methods demonstrate the potential of pruning to enable LLM deployment on resource-constrained devices without causing substantial performance degradation.
Existing pruning methods mostly adopt a layer-wise approach to deal with large-scale parameters in LLMs. Layer-wise pruning is a uniform sparsity strategy, applying the same pruning ratio across all layers of an LLM. While simple to implement, this approach overlooks the inherent differences in the contributions of different layers to the overall performance of the model. Thus, it can only find the local optimal pruning solution, but not the global optimal solution. Empirical evidence also suggests that certain layers are more critical than others, and uniformly pruning across all layers may lead to the removal of essential parameters, ultimately impairing the pruned model’s performance Mocanu et al. (2018); Yin et al. (2023).
To address this limitation, we leverage the concept of Shapley value from cooperative game theory and propose a novel Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs, which was originally designed to fairly assess the contributions of multiple players in a game. In this study, we treat each Transformer in an LLM as a “player”. Based on this idea, we evaluate the contribution of each Transformer to the overall performance of an LLM. The results are further used as a basis to allocate customized sparsity ratios to each Transformer, prioritizing the preservation of parameters in more important ones. This approach not only improves the performance of the pruned LLMs, but also introduces a theoretical foundation for sparsity allocation, moving beyond heuristic methods. We further propose a Sliding Window-based Shapley Value approximation method (SWSV) to reduce the computational cost.
We conduct extensive experiments to evaluate the performance of SV-NUP for pruning LLMs across a spectrum of LLMs, including LLaMA-v1, LLaMA-v2 and OPT. SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity. The results validate our hypothesis that non-uniform sparsity allocation can significantly enhance the performance of pruned models, paving the way for more efficient and practical LLM deployment on resource-constrained devices.
2 Related Work
Uniform Pruning. Traditional pruning requires a round of re-training to restore performance, which poses significant challenges for LLMs. Researchers have developed pruning algorithms specifically tailored for LLM compression. For instance, Ma et al. (2023) investigated structured sparse LLMs by applying Taylor pruning to remove entire weight rows, followed by LoRA fine-tuning Hu et al. (2021). In recent years, the focus has shifted toward unstructured pruning which eliminates the need for fine-tuning. SparseGPT Frantar and Alistarh (2023) employs the Hessian inverse for pruning, followed by weight updates to reduce reconstruction errors between dense and sparse weights. Wanda Sun et al. (2023) introduced a criterion that incorporates weight magnitude and input activations to preserve outlier features.
Non-uniform Pruning. Uniform layerwise sparsity is commonly used for pruning language models Zhu and Gupta (2017); Gale et al. (2019), with several studies demonstrating its effectiveness in LLM pruning Sanh et al. (2020); Kurtic et al. (2022). However, there is a growing body of work exploring non-uniform layerwise sparsity, primarily in the context of vision models. For example, Mocanu et al. (2016) proposed a non-uniform, scale-free topology inspired by graph theory, which outperforms dense counterparts when applied to restricted Boltzmann machines. Subsequent work has improved the scalability of this approach by leveraging Erdős–Rényi graphs ERDdS and R&wi (1959), extending the method to fully connected layers Mocanu et al. (2018) and convolutional layers Evci et al. (2020); Liu et al. (2022) to achieve data-free and feedforward-free layerwise sparsity. Another approach to non-uniform sparsity involves applying a global threshold across all layers Frankle and Carbin (2018); Lee et al. (2018); Wang et al. (2020); Liu et al. (2021). However, global pruning has been found to be computationally expensive and ineffective when applied to LLMs.
Analyzing LLMs. The authors in Shim et al. (2022) analyzed the contributions of various components in LLMs and their impact on overall performance. Gromov et al. (2024) explored the role of deep layers in LLMs through layer pruning, providing insights into how different layers in relation to model performance. Michel et al. (2019) examined the redundancy of attention heads in transformer-based models, demonstrating that many attention heads can be pruned without significant performance degradation. Clark (2019) investigated the behavior of individual attention heads in BERT, revealing that each head serves a distinct role in capturing different linguistic features. Probing techniques are widely used to analyze the internal representations of LLMs. For instance, Tenney (2019) employed probing tasks to examine the linguistic information captured by BERT, finding that different layers encode distinct types of linguistic features. Furthermore, Tenney et al. (2019) introduced a suite of probes to analyze the representations learned by contextualized word embeddings, offering insights into how syntactic and semantic information is distributed across layers. Recently, Zhang et al. (2024) used Shapley values to evaluate the importance of layers in LLMs, providing a faithful assessment of their contributions. However, these studies have not explored Transformer SV-based non-uniform pruning of LLMs. SV-NUP bridges this gap.
3 Motivation
In this section, we conduct empirical studies on three LLMs to demonstrate our motivation for allocating different pruning ratios for different Transformers. We select 3 comment LLMs: 1) TinyLLama, 2) LLama-7B and 3) Mistral-7B. We compare the difference between Transformers by mean and standard deviation (std) of the Magnitude method Lee et al. (2020), which is widely used in neural network pruning. Based on the results shown in Figure 1, we can make the following two observations.
3.1 Observations
(a) TinyLLaMA
(b) TinyLLaMA
(c) LLaMA-7B
(d) LLaMA-7B
(e) Mistral-7B
(f) Mistral-7B
Figure 1: Comparison of the mean and standard deviation (std) across different Transformers under Magnitude for 3 LLMs (TinyLLama, LLama-7B, and Mistral-7B). TinyLLama consists of 22 Transformers and the other 2 LLMs consist of the same 32 Transformers, where each Transformer consists of 7 innerlayers.
Observation 1
Different Transformers in LLMs exhibit significant variations in the mean and std of magnitude.
As the Transformer index increases, the mean magnitude decreases across all LLMs (Figures 1(a), 1(c) and 1(e)). Notably, the changes in LLaMA-7B and Mistral-7B are more pronounced compared to TinyLLaMa. The earlier (shallower) Transformers tend to have higher mean magnitudes. From the std of magnitude in Figures 1(b), 1(d) and 1(f), it can be observed that weights of shallow Transformers change dramatically. It is evident that the weights of shallow Transformers vary more significantly. Higher mean and std of magnitude values indicate that the correspondong Transformers have greater impact on LLM performance. These results underscore that different layers contribute unevenly to the overall model performance, suggesting that some Transformers are more sensitive to changes in parameter magnitudes.
Observation 2
The attention module of Transformers has a higher mean and std of magnitude than the feed-forward network (FFN) module.
Each Transformer consists of 7 inner layers. The first 4 correspond to the attention module, while the last 3 form the FFN module. In all the figures, the mean and std of magnitude in the attention module are higher than those in the FFN module. This finding indicates that the attention module is more sensitive compared to the FFN module to changes in parameter magnitude. For instance, in Figures 1(a) and 1(b), TinyLLaMa’s inner layer 4 shows a distinct dividing line: layers 1–4 form peaks, whereas layers 5–7 form troughs. The other three LLMs follow similar trends. As a result, the attention module warrants more attention during pruning.
3.2 Analysis and Implications
Conventional LLM pruning approaches that apply a uniform pruning ratio across all Transformers tend to overlook the aforementioned inherent diversity. For example, in Figure 1(c), the first few Transformers exhibit lower mean magnitudes, suggesting that they can withstand more aggressive pruning; whereas the latter Transformers show higher mean magnitudes and std, indicating that they should be pruned more conservatively to maintain performance. Applying the same pruning ratio to Transformers with high mean magnitudes or high std as those with lower values might remove critical capabilities from important Transformers, while preserving too many parameters in less significant ones.
This analysis clearly demonstrates the advantage of allocating different pruning ratios to different Transformers. By tailoring the pruning ratios based on the unique characteristics of each Transformer, pruning strategies can be optimized to retain crucial information in the model. Such a non-uniform pruning approach enables more efficient compression where less critical Transformers undergo more aggressive pruning, while vital Transformers are preserved, thereby enhancing the performance of the pruned model.
4 The Proposed SV-NUP Method
Motivated by these observations, we investigate the issue of adaptive sparsity for LLM pruning. The Shapley value, a concept derived from cooperative game theory, is widely used to fairly allocate contributions among multiple players in a game. This paper leverages the Shapley value to allocate pruning ratios across different Transformers. By treating each Transformer as a “player” in the model, we can compute the contribution of each Transformer to the overall performance of the LLM. This approach will allow us to assign pruning ratios dynamically, ensuring that more important Transformers (i.e., those with higher Shapley values) are pruned less aggressively, while less critical Transformers receive higher pruning ratios. We propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method to address this problem, which provides a theoretically grounded and equitable solution to the problem of non-uniform pruning. This method optimizes the trade-off between model size and performance, ensuring that critical Transformers are preserved, while less significant ones are pruned more aggressively.
We consider post-training pruning of LLMs from a well-optimized model with weights $\mathbf{W}^{*}$ to a sparse version $\mathbf{W}$ with many 0 values under a given pruning ratio $\rho$. As LLMs contain hundreds of millions of parameters, we usually adopt layer-wise pruning. The neural network structure of LLMs is generally composed of many Transformers. We denote Transformers as the set $\mathcal{T}=\{1,2,...,T\}$. Conventional LLM pruning methods apply a uniform pruning ratio $\rho$ to all Transformers. According to the analysis in Section 3, we should allocate different pruning ratios $\rho_{t}$ to different Transformer $t\in\mathcal{T}$. The key to allocating pruning ratios is to calculate the contribution of Transformers of LLMs precisely. We use SV to analyze each Transformer contribution.
4.1 Transformer Contribution by Shapley Value
SV is a unique contribution allocation scheme that satisfies a set of fairness axioms. Here, we leverage the concept to calculate the weighted average of all marginal contributions of each Transformer in an LLM. To illustrate the calculation process, we consider a simple LLM with 3 Transformers as shown in Figure 2. Since all Transformers have the same structure, we can easily mask a Transformer during inference and do not affect SV calculation, as shown in the bottom right corner of Figure 2. We use the accuracy of an LLM as the basis for calculating the SVs.
Figure 2: Contribution analysis of 3 Transformers in an LLM. For intuitive presentation, we use accuracy to represent the contribution of different Transformers.
We first identify each Transformer’s contribution when they participate individually, when 2 participate together, and when all 3 participate together. Particularly, as an LLM without any Transformer cannot function normally, we assume the accuracy to be 0. Then, we consider all possible combinations of Transformers and calculate their marginal values (e.g. what value does each Transformer add when Transformer 1 enters the LLM first, followed by Transformer 2, and then Transformer 3). Finally, we need to add them up and work out the SV (i.e. the average) for each Transformer.
To formalize this process, an LLM consists of $\mathcal{T}$ Transformers and a value function $\nu$. Perplexity (PPL) is usually used to evaluate LLMs in LLM pruning JAISWAL et al. (2024b); Xia et al. (2024); Ma et al. (2023). The lower the value, the better the performance. Therefore, the value function is denoted as $\nu(\mathcal{T})=1/\text{PPL}(\mathcal{T})$. Particularly, if there is no Transformer, the value function must be 0, i.e., $\nu(\emptyset)=0$. Let $\mathcal{S}\subseteq\mathcal{T}$ denote a subset of $\mathcal{T}$. According to Shapley (1953), the SV of any Transformer $t\in\mathcal{T}$ is:
$\phi_{t}=\sum_{\mathcal{S}\subseteq\mathcal{T}\setminus\{t\}}w_{\mathcal{S}}[\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})],\forall t\in\mathcal{T},$ (1)
where $w_{\mathcal{S}}=T\cdot\binom{T-1}{\mathcal{S}}=\frac{T!}{|\mathcal{S}|!(T-|\mathcal{S}|-1)!}$ is a coefficient. $\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})$ is known as Transformer $t$’s marginal contribution. Eq. (1) yields the following desirable properties:
  • Efficiency: The sum of all Shapley values equals to the total contribution towards an LLM by all Transformers, i.e., $\sum_{t\in\mathcal{T}}\phi_{t}=v(\mathcal{T})$. This ensures that all contributions from Transformers are fully allocated with no “leftover” or “missing” gain.
  • Symmetry: If two Transformers contribute equally to an LLM (i.e., they have the same marginal contribution in all subsets), they should have the same Shapley value, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}\cup\{i\}),\forall\mathcal{S}\subseteq\mathcal{T}\setminus\{t,i\}\Rightarrow\phi_{t}=\phi_{i}$. This ensures fairness, as Transformers with the same effect on an LLM are not favored over one another.
  • Dummy: If a Transformer does not add any marginal contribution to any subset of Transformers, its Shapley value is 0, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}),\forall\mathcal{S}\subseteq\mathcal{T}\Rightarrow\phi_{t}=0$. This ensures that Transformers that do not affect an LLM’s output receive no contribution value.
Eq. (1) involves calculating the marginal contribution for every subset $\mathcal{S}$, which results in a combinatorial explosion as the number of Transformers increases. Specifically, for $T$ Transformers, the total number of possible subsets $\mathcal{S}$ is $2^{T}$. For instance, with 32 Transformers in LLaMa-7B, the number of possible subsets becomes $2962^{32}=4,294,967,$ 296. The exact computation of Shapley values requires evaluating every subset that excludes the given Transformer, meaning that $32\times 2^{32}$ evaluations are needed. Each additional feature doubles the number of subsets that must be considered, causing an exponential increase in the number of computations. Consequently, this problem is NP-hard.
4.2 Shapley Value Approximation
Figure 3: Selection of Transformer coalitions under an LLM with 8 Transformers.
To address the NP-hard problem, this subsection proposes a sliding window-based Shapley value (SWSV) approximation method to estimate the contribution of Transformers. The key is that many coalitions of Transformers are useless and meaningless when evaluating LLMs. We consider an LLM with 8 Transformers as shown in Figure 3, where Transformers 1 and 8 are active and Transformers 2-7 are masked. We have two limitations: 1) too few active Transformers; and 2) too far apart between active Transformers.
Specifically, a well-trained LLM relies on maintaining a sufficiently large network structure, even in the presence of sparsity, to preserve its representation capacity and functionality Li et al. (2022); Lee et al. (2024). Sparsity, while effective in reducing computational costs, must be implemented judiciously to ensure that the structural integrity and critical pathways of the model remain intact. An LLM with an insufficient number of active Transformer layers—such as configurations involving only two active layers (e.g., Transformers 1 and 8)—is highly likely to suffer from significant performance degradation. This occurs because such configurations disrupt the hierarchical processing of information, resulting in incomplete or suboptimal feature representations. Moreover, when many intermediate Transformers are masked, the direct transmission of outputs from an earlier Transformer (e.g., Transformer 1) to a much later one (e.g., Transformer 8) bypasses the essential intermediate processing steps. This shortcut diminishes the model’s ability to learn and propagate nuanced information through the intermediate layers, leading to a distorted representation of the Shapley value. Such scenarios are problematic as they fail to reflect the contributions of individual layers within the model architecture.
To address these issues, it is critical to design an SV approximation method to avoid extreme patterns with few Transformers. The goal is to improve the accuracy of SV estimation, enhance model performance, and reduce computational complexity by preserving critical Transformers.
Figure 4: Illustration of the proposed approximation.
To this end, we propose a sliding window-based Shapley value approximation method for evaluating the contribution of Transformers in LLMs (Figure 4). When calculating the SV of Transformer $t\in\mathcal{T}$, we consider the sliding window with $N=|\mathcal{S}_{t}|$ Transformers, where these Transformers are closely connected. Figure 4 shows the sliding window with 3 Transformers, where the SV of Transformer 1 is only related to the Transformers in this sliding window. It is expressed as:
$\hat{\phi}_{t}=\sum_{\mathcal{S}\subseteq\mathcal{S}_{t}\setminus\{t\}}w_{\mathcal{S}}[\nu(\mathcal{S}\cup\mathcal{T}_{t}\cup\{t\})-\nu(\mathcal{S}\cup\mathcal{T}_{t})],\forall t\in\mathcal{T},$ (2)
where $\mathcal{T}_{t}=\mathcal{T}\setminus\mathcal{S}_{t}$ is coalitions without $\mathcal{S}_{t}$.
Neighboring Transformers around a target Transformer $t$ play a significant role in determining its SV, as they strongly influence the context and interactions that contribute to the target Transformer’s functionality. To effectively capture these interdependencies, a sliding window strategy is employed, which selects a set of neighboring Transformers both before and after the target Transformer. This approach ensures that the SV calculation takes into account the most relevant local interactions within the Transformer stack. The sliding window size $N$ is generally set to an odd number, such as 3, 5, or 7, to maintain symmetry around the target Transformer. For instance, with a sliding window of size 3, the calculation includes one preceding Transformer ($t-1$), the target Transformer itself ($t$), and one succeeding Transformer ($t+1$). This balanced selection ensures that contributions from both upstream and downstream Transformers are equally considered, facilitating a more accurate evaluation of the target Transformer’s importance.
Input: $\mathcal{T}$, $N$, an empty set $\mathcal{V}$, and a dataset;
Output: Shapley values of Transformers;
1
2while $t\in\mathcal{T}$ do
3       Determine the sliding window $\mathcal{S}_{t}$;
4       for $\mathcal{S}$ in all coalitions of $\mathcal{S}_{t}\setminus\{t\}$ do
5             if $\mathcal{S}\cup\mathcal{T}_{t}\cup\{t\}\not\subseteq\mathcal{V}$ then
6                   Put $\mathcal{S}\cup\mathcal{T}_{t}\cup\{t\}$ into $\mathcal{V}$;
7                   Calculate $\nu(\mathcal{S}\cup\mathcal{T}_{t}\cup\{t\})$;
8
9             end if
10            if $\mathcal{S}\cup\mathcal{T}_{t}\not\subseteq\mathcal{V}$ then
11                   Put $\mathcal{S}\cup\mathcal{T}_{t}$ into $\mathcal{V}$;
12                   Calculate $\nu(\mathcal{S}\cup\mathcal{T}_{t})$;
13
14             end if
15
16       end for
17      Calculate $\hat{\phi}_{t}$ by (2);
18
19 end while
Algorithm 1 SWSV
Special considerations are made for boundary Transformers, such as the first and last ones, where the default sliding window size might result in insufficient neighbors. In these cases, the sliding window is adjusted to include the available Transformers, while maintaining the specified number of neighbors as much as possible. For example, when evaluating the first Transformer, additional downstream Transformers can be included to compensate for the absence of preceding Transformers. Similarly, for the last Transformer, upstream Transformers are prioritized. This ensures consistency in the Shapley value calculations across all Transformers, preserving the intended neighborhood size while respecting the structural boundaries of the model. The pseudocode of SWSV is shown in Algorithm 1. Its computational complexity is $\mathcal{O}(T2^{N-1})$, which is much lower than the original complexity.
4.3 SWSV for Pruning LLMs
After calculating the contributions of Transformers (${\hat{\phi}_{t}},{t\in\mathcal{T}}$), these contributions can be utilized to allocate the pruning ratio for each Transformer. Transformers with greater contributions are considered more important and thus have smaller pruning ratios. Therefore, the pruning ratio $\rho_{t}$ is inversely proportional to the contribution, i.e., $\rho_{t}\propto 1/\hat{\phi}_{t}$. However, since overly sparse Transformers can degrade the performance of LLMs, it is necessary to constrain the allocated pruning ratios. To achieve this, we introduce a small positive value $\lambda$ (e.g., 0.05 or 0.1) as a boundary. The allocated pruning ratios are restricted to lie within the given pruning ratio range, with an allowable error of $\lambda$, i.e., $\rho_{t}\in[\rho-\lambda,\rho+\lambda]$. The detailed mathematical expression is:
$\rho_{t}=\rho-a_{t}+\text{mean}\{a_{t}\}_{t\in\mathcal{T}},$ (3)
where $a_{t}=\frac{2\lambda(\hat{\phi}_{t}-\min{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}})}{\max{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}}-\min{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}}}$.
SV-NUP is a framework which can be applied to advanced pruning methods (e.g., Wanda Sun et al. (2023) and SparseGPT Frantar and Alistarh (2023)). The pseudocode of SV-NUP is presented in Algorithm 2.
Input: an LLM and a dataset;
Output: The Pruned LLM;
1 Calculate the contributions $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$ of Transformers by SWSV;
2 Allocate the pruning ratios $\{\rho_{t}\}_{t\in\mathcal{T}}$ by Eq. (3);
3 while $t\in\mathcal{T}$ do
4       Prune the Transformer $t$ by an advanced pruning method at the pruning ratio $\rho_{t}$;
5
6 end while
7Evaluate the pruned LLM;
Algorithm 2 SV-NUP
5 Experimental Evalution
5.1 Experiment Settings
We evaluate the proposed SV-NUP method on the LLaMA-V1, LLaMA-V2, and OPT model families Frantar and Alistarh (2023). Due to computational resource constraints, our focus is on the 7-billion and 13-billion parameter variants. All LLMs are quantized with float16 and the sequence length is set as 2048. To assess the performance of the pruned LLMs, we employ two general metrics: PPL and zero-shot evaluation Sun et al. (2023). PPL reflects the model’s ability to predict the next word given the preceding context Muhlgay et al. (2023), where lower values indicate better performance. We evaluate PPL using the WikiText-2 dataset Merity et al. (2016), selecting two randomly downloaded subsets of WikiText-2 for testing. For zero-shot evaluation, we choose 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness Mihaylov et al. (2018), where higher scores indicate better performance.
Table 1: PPL↓  performance of LLaMA family pruned by 3 existing methods with/without (w./wo.) SV-NUP at 70% sparsity. We highlight the improved performance in blue.
LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B
Dense 5.68 5.09 5.47 4.88
Magnitude 48426.42 84531.15 49808.59 214.21
SV-NUP w. Magnitude 30586.90 (-36.84%) 10632.55 (-87.42%) 31729.74 (-36.30%) 82.98 (-61.26%)
Wanda 70.84 41.10 57.20 34.06
SV-NUP w. Wanda 34.95 (-50.66%) 15.92 (-61.28%) 29.56 (-48.33%) 20.46 (-39.94%)
SparseGPT 18.42 13.74 17.87 14.39
SV-NUP w. SparseGPT 15.10 (-18.01%) 11.06 (-19.55%) 14.84 (-17.00%) 12.48 (-13.27%)
5.2 Comparison Baselines
We select 3 baselines for comparison: 1) Magnitude Jaiswal et al. (2024a), 2) Wanda Sun et al. (2023), and 3) SparseGPT Frantar and Alistarh (2023). The proposed SV-NUP method is integrated into these baselines, allowing us to directly observe the performance improvement brought by it. In addition, we compare SV-NUP with OWL Yin et al. (2023), a pruning ratio allocation framework based on outliers.
We implement all experiments in PyTroch and use the HuggingFace library to download the pre-trained LLMs and datasets. All experiments are performed on a server with 4 NVIDIA A100 GPUs, 1 AMD EPYC CPU, and 252 GB of memory. The limitation $\lambda$ of the pruning ratio is set to 0.1. The size of sliding windows $N$ is set as 3, 5, or 7.
5.3 PPL Improvement by SV-NUP
We first evaluate the PPL performance gains achieved by integrating SV-NUP into LLM pruning methods, as shown in Table 1. SV-NUP significantly enhances existing pruning approaches, particularly Magnitude and Wanda, while SparseGPT demonstrates minimal improvements due to its already strong baseline performance. SV-NUP has 61.28 % improvement on LLaMA-13B by Wanda. Wanda consistently benefits from SV-NUP across all LLaMA models, with the largest impact observed in larger models like LLaMA-13B, where SV-NUP effectively mitigates performance degradation. SparseGPT, despite its high baseline effectiveness, achieves further refinement with SV-NUP, highlighting the compatibility and synergy between the two. Overall, SV-NUP emerges as a crucial enhancement, stabilizing pruning outcomes and delivering superior results, especially when combined with advanced methods like SparseGPT.
Table 2: PPL↓  performance of LLaMA family and OPT family pruned by different methods at 50% sparsity. We highlight the best performance in bold.
TinyLLaMA LLaMA-7B Vicuna-7B LLaMA2-7B LLaMA3.2-3B OPT-2.7B OPT-6.7B
Dense 7.7735 5.6772 6.9031 5.4721 7.8137 12.4705 10.8602
Magnitude 21.6367 17.2882 24.0034 16.0301 139.4124 265.2033 968.7209
OWL w. Magnitude 18.7673 16.3453 21.3578 15.7421 99.2163 207.0598 363.6955
SV-NUP w. Magnitude 18.1645 15.9755 20.7409 15.3144 89.4710 158.2674 308.4234
Wanda 10.8549 7.0884 8.5325 6.7741 12.6977 13.9611 12.0780
OWL w. Wanda 11.1533 7.0941 8.5731 6.7954 12.6735 14.7622 12.3921
SV-NUP w. Wanda 11.0834 7.0610 8.4428 6.7959 12.1578 14.0323 12.0065
SparseGPT 9.8810 6.8701 8.0767 6.6000 11.0357 12.0591 11.2711
OWL w. SparseGPT 10.1521 6.8785 8.2247 6.6388 10.9828 13.3490 11.5101
SV-NUP w. SparseGPT 10.1308 6.8336 8.0762 6.5843 10.8495 12.9768 11.2582
Table 3: PPL↓  and classification performance of Vicuna-7B and OPT-6.7B pruned by different methods at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
PPL↓ BoolQ RTE WG OBQA Mean
Vicuna-7B Dense 6.9031 0.7810 0.6823 0.6938 0.3460 0.6258
Magnitude 24.0034 0.5462 0.5271 0.5817 0.2200 0.4687
OWL w. Magnitude 21.3578 0.6067 0.5784 0.6056 0.2300 0.5052
SV-NUP w. Magnitude 20.7409 0.6037 0.5884 0.5959 0.2360 0.5060
Wanda 8.5325 0.6969 0.6859 0.6519 0.2960 0.5827
OWL w. Wanda 8.5731 0.6315 0.7040 0.6685 0.2960 0.5750
SV-NUP w. Wanda 8.4428 0.6976 0.6968 0.6630 0.2880 0.5863
SparseGPT 8.0767 0.6437 0.6534 0.6480 0.2940 0.5598
OWL w. SparseGPT 8.2247 0.6780 0.6715 0.6472 0.2840 0.5702
SV-NUP w. SparseGPT 8.0762 0.6945 0.6823 0.6567 0.3020 0.5839
OPT-6.7B Dense 10.8602 0.6606 0.5523 0.6519 0.2760 0.5352
Magnitude 968.7209 0.3804 0.5271 0.5059 0.1760 0.3974
OWL w. Magnitude 363.6955 0.3813 0.5271 0.5564 0.2140 0.4197
SV-NUP w. Magnitude 308.4234 0.3908 0.5271 0.5567 0.2120 0.4217
Wanda 12.0780 0.6214 0.5271 0.6014 0.2380 0.4970
OWL w. Wanda 12.3921 0.6220 0.5271 0.5864 0.2380 0.4934
SV-NUP w. Wanda 12.0065 0.6211 0.5271 0.6069 0.2420 0.4993
SparseGPT 11.2711 0.6358 0.5343 0.6433 0.2540 0.5168
OWL w. SparseGPT 11.5101 0.6324 0.5307 0.6393 0.2320 0.5086
SV-NUP w. SparseGPT 11.2582 0.6410 0.5543 0.6338 0.2540 0.5208
Table 4: PPL↓  performance under different $N$ at 50% sparsity.
LLaMA-7B $N=3$ $N=5$ $N=7$
SV-NUP w. Magnitude 16.0491 16.0081 15.9755
SV-NUP w. Wanda 7.09322 7.06104 7.09215
SV-NUP w. SparseGPT 6.8382 6.83358 6.84875
OPT-6.7B $N=3$ $N=5$ $N=7$
SV-NUP w. Magnitude 525.407 308.423 661.884
SV-NUP w. Wanda 12.0469 12.0339 12.0065
SV-NUP w. SparseGPT 11.2766 11.2661 11.2582
Table 5: Pruning ratio allocation by two SV approximation methods on LLaMA-7B and OPT-6.7B at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
PPL↓ BoolQ RTE WG OBQA Mean
LLaMA-7B SV-NUP w. Magnitude 15.98 0.58 0.58 0.61 0.27 0.5100
SV w. Magnitude 17.48 0.56 0.57 0.61 0.25 0.4956
SV-NUP w. Wanda 7.06 0.73 0.59 0.68 0.30 0.5638
SV w. Wanda 7.11 0.73 0.55 0.67 0.30 0.5626
SV-NUP w. SparseGPT 6.83 0.73 0.55 0.69 0.27 0.5584
SV w. SparseGPT 6.86 0.72 0.52 0.68 0.28 0.5493
OPT-6.7B SV-NUP w. Magnitude 308.42 0.39 0.53 0.56 0.21 0.4217
SV w. Magnitude 660.89 0.38 0.53 0.53 0.20 0.4089
SV-NUP w. Wanda 12.01 0.62 0.53 0.61 0.24 0.4993
SV w. Wanda 12.06 0.62 0.53 0.61 0.23 0.4977
SV-NUP w. SparseGPT 11.26 0.64 0.55 0.63 0.25 0.5208
SV w. SparseGPT 11.33 0.63 0.53 0.63 0.25 0.5111
Figure 5: Distribution of the allocated pruning ratio on LLaMA-7B by 3 methods at 50% sparsity.
5.4 PPL and Zero-shot Performance by Baselines
The results in Table 2 underscore the effectiveness of integrating SV-NUP and OWL into existing pruning methods for the LLaMA and OPT model families at 50% sparsity. Across most scenarios, SV-NUP consistently demonstrates superior performance. Specifically, Magnitude combined with SV-NUP consistently outperforms both its standalone version and Magnitude with OWL. Notably, while SV-NUP slightly improves the performance of Wanda and SparseGPT, the gains are less pronounced, as the pruned LLMs at 50% sparsity already exhibit robust performance, leaving limited room for further enhancements. Interestingly, SV-NUP’s effectiveness is more apparent with larger LLMs, such as the 7-billion-parameter models, whereas its impact diminishes for smaller models like TinyLLaMA and OPT-2.7B. For Wanda, the integration of SV-NUP results in a notable reduction in PPL compared to Wanda alone, highlighting the synergy between SV-NUP and fine-grained pruning strategies. OWL also contributes positively, particularly when integrated with SparseGPT and Wanda, though its impact is slightly less consistent than that of SV-NUP. Overall, integrating SV-NUP stabilizes and enhances performance across various pruning methods, emerging as a particularly strong addition. It consistently ensures competitive results and maintains lower PPL across a wide range of LLM architectures.
Table 3 presents the PPL and classification performance of OPT-2.7B and OPT-6.7B models pruned by different methods at 50% sparsity. For OPT-2.7B, SV-NUP with SparseGPT achieves a mean score of 0.4951, which is slightly lower than SparseGPT alone (0.5039), suggesting that the integration of SV-NUP with SparseGPT may not provide performance gains and could slightly degrade results for smaller OPT models. In contrast, for OPT-6.7B, SV-NUP with SparseGPT achieves a mean score of 0.5208, which is comparable to SparseGPT alone (0.5168), indicating that SV-NUP can maintain performance levels when combined with SparseGPT for larger models.
Table 4 illustrates the impact of varying sliding window sizes ($N$) on performance. The results indicate that $N=5$ and $N=7$ achieve lower PPL for both LLaMA-7B and OPT-6.7B. However, the choice of $N$ is highly sensitive to different LLMs and pruning methods, highlighting the importance of selecting an appropriate window size tailored to specific scenarios.
5.5 Different SV Approximation Methods
We compare our proposed SWSV method with the existing SV approximation method proposed in Kolpaczki et al. (2024) to demonstrate its efficiency, as shown in Table 5. For LLaMA-7B, SWSV with SparseGPT achieves the highest mean score (0.5584), slightly surpassing SV with SparseGPT (0.5493), showcasing its ability to enhance performance. Similarly, for OPT-6.7B, SWSV with SparseGPT achieves a mean score of 0.5208, outperforming SV with SparseGPT (0.5111). These results demonstrate that SWSV not only provides marginal performance improvements over standard SV, particularly when combined with SparseGPT, but also offers lower computational complexity. This highlights its potential as a more efficient and effective sparsity allocation approach.
Figure 5 shows the distribution of allocated pruning ratios for LLaMA-7B at 50% sparsity across three methods: SWSV, SV, and OWL. SWSV demonstrates significant variability, particularly in earlier layers, with pruning ratios fluctuating more widely before stabilizing closer to the latter layers.
6 Conclusions and Future Work
In this paper, we propose SV-NUP, a novel approach for non-uniform pruning of LLMs based on the contribution by each Transformer to overall model performance. By leveraging the Shapley value, we propose a method to assess the importance of individual Transformers within LLMs. To further address the computational complexity associated with the Shapley value, we design a sliding window-based approximation method. Extensive experiments have been carried out on LLMs including LLaMA-V1, LLaMA-V2 and OPT model families in comparison with three state-of-the-art pruning methods. SV-NUP achieves significant improvements in both PPL and zero-shot performance, demonstrating its promise as a useful LLM pruning method that can better preserve model performance at the given pruning ratios.
In subsequent research, we plan to investigate the integration of pruning and quantization for compressing LLMs to reduce computation and storage.

### v3

@@
 Pruning large language models (LLMs) is a promising solution for reducing model sizes and computational complexity while preserving performance.
-Traditional layer-wise pruning methods often adopt a uniform sparsity approach across all layers, which leads to suboptimal performance due to the varying significance of individual transformer layers within the model not being accounted for.
+Traditional layer-wise pruning methods often adopt uniform sparsity at all Transformer layers, which leads to suboptimal performance due to the varying significance of transformer layers not being accounted for.
 To this end, we propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs.
-This approach quantifies the contribution of each transformer layer to the overall model performance, enabling the assignment of tailored pruning budgets to different layers to retain critical parameters.
-To further improve efficiency, we design the Sliding Window-based Shapley Value approximation method.
+This method quantifies the contribution of each transformer layer to the overall model performance, enabling the assignment of tailored pruning budgets to different layers to retain critical parameters.
+To further improve efficiency, we design the Sliding Window-based Shapley Value (SWSV) approximation method.
 It substantially reduces computational overhead compared to exact SV calculation methods.
-Extensive experiments on various LLMs including LLaMA-v1, LLaMA-v2 and OPT demonstrate the effectiveness of the proposed approach.
+Extensive experiments on various LLMs including LLaMA-v1/v2/v3, and OPT demonstrate the effectiveness of SV-NUP.
 The results reveal that non-uniform pruning significantly enhances the performance of pruned models.
@@
 1 Introduction
-Large language models (LLMs) have emerged as a transformative technology, demonstrating remarkable capabilities in natural language understanding and generation tasks Brown et al.
-(2020).
+Large language models (LLMs) have emerged as a transformative technology, demonstrating remarkable capabilities in natural language understanding and generation tasks [1].
 LLMs have showcased significant adaptability through fine-tuning, enabling their deployment in highly specialized applications.
-These advantages underscore the critical role of LLMs in solving real-world challenges Patterson et al.
-(2021).
+These advantages underscore the critical role of LLMs in solving real-world challenges [2].
 Despite these capabilities, the deployment of LLMs has been hindered by their immense computational demands.
-Modern LLMs often consist of billions or even trillions of parameters Luccioni et al.
-(2023), as seen in models like GPT-3 (175 billion parameters) and PaLM (540 billion parameters).
+Modern LLMs often consist of billions or even trillions of parameters [3], as seen in models like GPT-3 (175 billion parameters) and PaLM (540 billion parameters).
 The immense scale incurs significant memory, storage and power costs, making it challenging to run these models on resource-constrained devices.
-To address these issues, researchers have increasingly turned to model compression techniques, particularly pruning, to reduce the size and computational requirements of LLMs, while retaining their performance Frantar and Alistarh (2023).
-Pruning LeCun et al.
-(1989); Han et al.
-(2015b, a) is one of the most prominent model compression techniques.
+To address these issues, researchers have increasingly turned to model compression techniques, particularly pruning, to reduce the size and computational requirements of LLMs, while retaining their performance [4].
+Pruning [5; 6; 7] is one of the most prominent model compression techniques.
 It aims to remove redundant or less important parameters from neural networks, thereby reducing their sizes and computational complexity.
-Recent advances in pruning have progressed from unstructured sparsity (i.e., individual weight removal) to structured sparsity (i.e., eliminating entire neurons, heads or layers), thereby enabling hardware-efficient implementation Frantar and Alistarh (2023); Jaiswal et al.
-(2024a); Ma et al.
-(2023); Xia et al.
-(2023).
+Recent advances in pruning have progressed from unstructured sparsity (i.e., individual weight removal) to structured sparsity (i.e., eliminating entire neurons, heads or layers), thereby enabling hardware-efficient implementation [4; 8; 9; 10].
 For LLMs, several state-of-the-art methods have been proposed, including magnitude pruning, lottery ticket hypothesis approaches, and structured pruning based on attention mechanisms.
 These methods demonstrate the potential of pruning to enable LLM deployment on resource-constrained devices without causing substantial performance degradation.
-Existing pruning methods mostly adopt a layer-wise approach to deal with large-scale parameters in LLMs.
-Layer-wise pruning is a uniform sparsity strategy, applying the same pruning ratio across all layers of an LLM.
+Figure 1: Overview and effectiveness of SV-NUP: (a) Distribution of weight magnitude across Transformer layers; (b) Conceptual comparison of uniform pruning vs.
+SV-NUP; (c) PPL↓  (the smaller the better) comparison under different pruning strategies.
+The red number is the improvement percentage.
+Unlike uniform pruning, which ignores layer-wise importance and may degrade LLM performance, SV-NUP leverages Shapley value to estimate the contribution of each Transformer layer and allocates pruning ratios accordingly to better preserve model performance.
+Existing pruning methods often use a layer-wise strategy that applies uniform sparsity across all layers, ignoring their varying importance.
 While simple to implement, this approach overlooks the inherent differences in the contributions of different layers to the overall performance of the model.
 Thus, it can only find the local optimal pruning solution, but not the global optimal solution.
-Empirical evidence also suggests that certain layers are more critical than others, and uniformly pruning across all layers may lead to the removal of essential parameters, ultimately impairing the pruned model’s performance Mocanu et al.
-(2018); Yin et al.
-(2023).
+Empirical evidence also suggests that certain layers are more critical than others, and uniformly pruning across all layers may lead to the removal of essential parameters, ultimately impairing the pruned model’s performance [11; 12].
 To address this limitation, we leverage the concept of Shapley value from cooperative game theory and propose a novel Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs, which was originally designed to fairly assess the contributions of multiple players in a game.
-In this study, we treat each Transformer in an LLM as a “player”.
-Based on this idea, we evaluate the contribution of each Transformer to the overall performance of an LLM.
-The results are further used as a basis to allocate customized sparsity ratios to each Transformer, prioritizing the preservation of parameters in more important ones.
+In this study, we treat each Transformer layer in an LLM as a “player".
+Based on this idea, we evaluate the contribution of each Transformer layer to the overall performance of an LLM.
+The results are further used as a basis to allocate customized sparsity ratios to each Transformer layer, prioritizing the preservation of parameters in more important ones.
 This approach not only improves the performance of the pruned LLMs, but also introduces a theoretical foundation for sparsity allocation, moving beyond heuristic methods.
 We further propose a Sliding Window-based Shapley Value approximation method (SWSV) to reduce the computational cost.
-We conduct extensive experiments to evaluate the performance of SV-NUP for pruning LLMs across a spectrum of LLMs, including LLaMA-v1, LLaMA-v2 and OPT.
+The overview of SV-NUP is shown in Figure 1.
+We conduct extensive experiments to evaluate the performance of SV-NUP for pruning LLMs across a spectrum of LLMs, including LLaMA-v1/v2/v3, and OPT.
 SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity.
@@
 2 Related Work
-Uniform Pruning.
-Traditional pruning requires a round of re-training to restore performance, which poses significant challenges for LLMs.
-Researchers have developed pruning algorithms specifically tailored for LLM compression.
-For instance, Ma et al.
-(2023) investigated structured sparse LLMs by applying Taylor pruning to remove entire weight rows, followed by LoRA fine-tuning Hu et al.
-(2021).
-In recent years, the focus has shifted toward unstructured pruning which eliminates the need for fine-tuning.
-SparseGPT Frantar and Alistarh (2023) employs the Hessian inverse for pruning, followed by weight updates to reduce reconstruction errors between dense and sparse weights.
-Wanda Sun et al.
-(2023) introduced a criterion that incorporates weight magnitude and input activations to preserve outlier features.
-Non-uniform Pruning.
-Uniform layerwise sparsity is commonly used for pruning language models Zhu and Gupta (2017); Gale et al.
-(2019), with several studies demonstrating its effectiveness in LLM pruning Sanh et al.
-(2020); Kurtic et al.
-(2022).
-However, there is a growing body of work exploring non-uniform layerwise sparsity, primarily in the context of vision models.
-For example, Mocanu et al.
-(2016) proposed a non-uniform, scale-free topology inspired by graph theory, which outperforms dense counterparts when applied to restricted Boltzmann machines.
-Subsequent work has improved the scalability of this approach by leveraging Erdős–Rényi graphs ERDdS and R&wi (1959), extending the method to fully connected layers Mocanu et al.
-(2018) and convolutional layers Evci et al.
-(2020); Liu et al.
-(2022) to achieve data-free and feedforward-free layerwise sparsity.
-Another approach to non-uniform sparsity involves applying a global threshold across all layers Frankle and Carbin (2018); Lee et al.
-(2018); Wang et al.
-(2020); Liu et al.
-(2021).
-However, global pruning has been found to be computationally expensive and ineffective when applied to LLMs.
-Analyzing LLMs.
-The authors in Shim et al.
-(2022) analyzed the contributions of various components in LLMs and their impact on overall performance.
-Gromov et al.
-(2024) explored the role of deep layers in LLMs through layer pruning, providing insights into how different layers in relation to model performance.
-Michel et al.
-(2019) examined the redundancy of attention heads in transformer-based models, demonstrating that many attention heads can be pruned without significant performance degradation.
-Clark (2019) investigated the behavior of individual attention heads in BERT, revealing that each head serves a distinct role in capturing different linguistic features.
-Probing techniques are widely used to analyze the internal representations of LLMs.
-For instance, Tenney (2019) employed probing tasks to examine the linguistic information captured by BERT, finding that different layers encode distinct types of linguistic features.
-Furthermore, Tenney et al.
-(2019) introduced a suite of probes to analyze the representations learned by contextualized word embeddings, offering insights into how syntactic and semantic information is distributed across layers.
-Recently, Zhang et al.
-(2024) used Shapley values to evaluate the importance of layers in LLMs, providing a faithful assessment of their contributions.
-However, these studies have not explored Transformer SV-based non-uniform pruning of LLMs.
-SV-NUP bridges this gap.
+LLM pruning often requires retraining to recover performance, which presents substantial challenges.
+While existing LLM-specific pruning methods primarily adopt uniform pruning strategies [9; 13; 4; 14], non-uniform layerwise sparsity has been extensively studied in vision models [15; 16; 11; 17; 18; 19; 20; 21; 22].
+However, global pruning is often inefficient and computationally expensive for LLMs.
+Although prior work has analyzed the importance of LLM components [23; 24; 25; 26; 27; 28; 29], none have explored Shapley value (SV)-based non-uniform pruning.
+SV-NUP fills this gap.
+For a more detailed literature review, refer to Appendix A.
 3 Motivation
-In this section, we conduct empirical studies on three LLMs to demonstrate our motivation for allocating different pruning ratios for different Transformers.
-We select 3 comment LLMs: 1) TinyLLama, 2) LLama-7B and 3) Mistral-7B.
-We compare the difference between Transformers by mean and standard deviation (std) of the Magnitude method Lee et al.
-(2020), which is widely used in neural network pruning.
-Based on the results shown in Figure 1, we can make the following two observations.
+In this section, we conduct empirical studies on three LLMs (TinyLLama, LLaMA-7B, and Mistral-7B) to motivate the need for allocating different pruning ratios across Transformer layers.
+We analyze the differences using the Magnitude method [30] and summarize each Transformer layer’s structure with the mean and standard deviation of innerlayer metrics.
+As shown in Figure 2, two key observations can be drawn from the results.
+(a) TinyLLaMA with Mean
+(b) LLaMA-7B with Mean
+(c) Mistral-7B with Mean
+(d) TinyLLaMA with Std
+(e) LLaMA-7B with Std
+(f) Mistral-7B with Std
+Figure 2: Comparison of the mean and standard deviation (Std) across different Transformer layers under Magnitude.
+TinyLLaMA consists of 22 Transformer layers, LLama-7B and Mistral-7B consist of the same 32 Transformer layers, where each Transformer layer consists of 7 inner-layers.
 3.1 Observations
-(a) TinyLLaMA
-(b) TinyLLaMA
-(c) LLaMA-7B
-(d) LLaMA-7B
-(e) Mistral-7B
-(f) Mistral-7B
-Figure 1: Comparison of the mean and standard deviation (std) across different Transformers under Magnitude for 3 LLMs (TinyLLama, LLama-7B, and Mistral-7B).
-TinyLLama consists of 22 Transformers and the other 2 LLMs consist of the same 32 Transformers, where each Transformer consists of 7 innerlayers.
-Observation 1
-Different Transformers in LLMs exhibit significant variations in the mean and std of magnitude.
-As the Transformer index increases, the mean magnitude decreases across all LLMs (Figures 1(a), 1(c) and 1(e)).
+Obervation 1 Different Transformer layers in LLMs exhibit significant variations in the mean and std of magnitude.
+As the Transformer index increases, the mean magnitude decreases across all LLMs (Figures 2(a), 2(b) and 2(c)).
 Notably, the changes in LLaMA-7B and Mistral-7B are more pronounced compared to TinyLLaMa.
-The earlier (shallower) Transformers tend to have higher mean magnitudes.
-From the std of magnitude in Figures 1(b), 1(d) and 1(f), it can be observed that weights of shallow Transformers change dramatically.
-It is evident that the weights of shallow Transformers vary more significantly.
-Higher mean and std of magnitude values indicate that the correspondong Transformers have greater impact on LLM performance.
-These results underscore that different layers contribute unevenly to the overall model performance, suggesting that some Transformers are more sensitive to changes in parameter magnitudes.
-Observation 2
-The attention module of Transformers has a higher mean and std of magnitude than the feed-forward network (FFN) module.
-Each Transformer consists of 7 inner layers.
+The earlier (shallower) Transformer layers tend to have higher mean magnitudes.
+From the std of magnitude in Figures 2(d), 2(e) and 2(f), it can be observed that weights of shallow Transformer layers change dramatically.
+It is evident that the weights of shallow Transformer layers vary more significantly.
+Higher mean and std of magnitude values indicate that the corresponding Transformer layers have a greater impact on LLM performance.
+These results underscore that different layers contribute unevenly to the overall model performance, suggesting that some Transformer layers are more sensitive to changes in parameter magnitudes.
+Obervation 2 The attention module of Transformer layers has a higher mean and std of magnitude than the feed-forward network (FFN) module.
+Each Transformer layer consists of 7 inner layers.
 The first 4 correspond to the attention module, while the last 3 form the FFN module.
@@
 This finding indicates that the attention module is more sensitive compared to the FFN module to changes in parameter magnitude.
-For instance, in Figures 1(a) and 1(b), TinyLLaMa’s inner layer 4 shows a distinct dividing line: layers 1–4 form peaks, whereas layers 5–7 form troughs.
+For instance, in Figures 2(a) and 2(d), TinyLLaMa’s inner layer 4 shows a distinct dividing line: layers 1–4 form peaks, whereas layers 5–7 form troughs.
 The other three LLMs follow similar trends.
@@
 3.2 Analysis and Implications
-Conventional LLM pruning approaches that apply a uniform pruning ratio across all Transformers tend to overlook the aforementioned inherent diversity.
-For example, in Figure 1(c), the first few Transformers exhibit lower mean magnitudes, suggesting that they can withstand more aggressive pruning; whereas the latter Transformers show higher mean magnitudes and std, indicating that they should be pruned more conservatively to maintain performance.
-Applying the same pruning ratio to Transformers with high mean magnitudes or high std as those with lower values might remove critical capabilities from important Transformers, while preserving too many parameters in less significant ones.
-This analysis clearly demonstrates the advantage of allocating different pruning ratios to different Transformers.
-By tailoring the pruning ratios based on the unique characteristics of each Transformer, pruning strategies can be optimized to retain crucial information in the model.
-Such a non-uniform pruning approach enables more efficient compression where less critical Transformers undergo more aggressive pruning, while vital Transformers are preserved, thereby enhancing the performance of the pruned model.
+Conventional LLM pruning approaches apply uniform pruning across all Transformer layers, ignoring differences across Transformer layers.
+As shown in Figure 2(b), early Transformer layers exhibit lower mean magnitudes, suggesting that they can withstand more aggressive pruning; whereas the latter Transformer layers show higher mean magnitudes and std, indicating that they should be pruned more conservatively to maintain performance.
+Applying the same pruning ratio to Transformer layers with high mean magnitudes or high std as those with lower values might remove critical capabilities from important Transformer layers, while preserving too many parameters in less significant ones.
+This analysis clearly demonstrates the advantage of allocating different pruning ratios to different Transformer layers.
+By tailoring the pruning ratios based on the unique characteristics of each Transformer layer, pruning strategies can be optimized to retain crucial information in the model.
+Such a non-uniform pruning approach enables more efficient compression where less critical Transformer layers undergo more aggressive pruning, while vital Transformer layers are preserved, thereby enhancing the performance of the pruned model.
 4 The Proposed SV-NUP Method
@@
 The Shapley value, a concept derived from cooperative game theory, is widely used to fairly allocate contributions among multiple players in a game.
-This paper leverages the Shapley value to allocate pruning ratios across different Transformers.
-By treating each Transformer as a “player” in the model, we can compute the contribution of each Transformer to the overall performance of the LLM.
-This approach will allow us to assign pruning ratios dynamically, ensuring that more important Transformers (i.e., those with higher Shapley values) are pruned less aggressively, while less critical Transformers receive higher pruning ratios.
-We propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method to address this problem, which provides a theoretically grounded and equitable solution to the problem of non-uniform pruning.
-This method optimizes the trade-off between model size and performance, ensuring that critical Transformers are preserved, while less significant ones are pruned more aggressively.
-We consider post-training pruning of LLMs from a well-optimized model with weights $\mathbf{W}^{*}$ to a sparse version $\mathbf{W}$ with many 0 values under a given pruning ratio $\rho$.
+This paper leverages the Shapley value to allocate pruning ratios across different Transformer layers.
+By treating each Transformer layer as a “player" in LLMs, we can compute the contribution of each Transformer layer to the overall performance of an LLM.
+This approach will allow us to assign pruning ratios dynamically, ensuring that more important Transformer layers (i.e., those with higher Shapley values) are pruned less aggressively, while less critical Transformer layers receive higher pruning ratios.
+We propose SV-NUP to address this problem, which provides a theoretically grounded and equitable solution to non-uniform pruning.
+This method optimizes the trade-off between model size and performance, ensuring that critical Transformer layers are preserved, while less significant ones are pruned more aggressively.
+We consider post-training pruning of LLMs from a well-optimized model with weights $\mathbf{W}^{*}$ to a sparse version $\mathbf{W}$ with many 0 under a given pruning ratio $\rho$.
 As LLMs contain hundreds of millions of parameters, we usually adopt layer-wise pruning.
-The neural network structure of LLMs is generally composed of many Transformers.
-We denote Transformers as the set $\mathcal{T}=\{1,2,...,T\}$.
-Conventional LLM pruning methods apply a uniform pruning ratio $\rho$ to all Transformers.
-According to the analysis in Section 3, we should allocate different pruning ratios $\rho_{t}$ to different Transformer $t\in\mathcal{T}$.
-The key to allocating pruning ratios is to calculate the contribution of Transformers of LLMs precisely.
-We use SV to analyze each Transformer contribution.
-4.1 Transformer Contribution by Shapley Value
+The neural network structure of LLMs is generally composed of many Transformer layers.
+We denote Transformer layers as the set $\mathcal{T}=\{1,2,...,T\}$.
+Conventional LLM pruning methods apply a uniform pruning ratio $\rho$ to all Transformer layers.
+According to the analysis in Section 3, we should allocate different pruning ratios $\rho_{t}$ to different Transformer layer $t\in\mathcal{T}$.
+The key to allocating pruning ratios is to calculate the contribution of Transformer layers of LLMs precisely.
+We use SV to analyze each Transformer layer contribution.
+4.1 Transformer Layer Contribution by Shapley Value
+Figure 3: Contribution analysis of 3 Transformer layers in an LLM. For intuitive presentation, we use accuracy to represent the contribution of different Transformer layers.
 SV is a unique contribution allocation scheme that satisfies a set of fairness axioms.
-Here, we leverage the concept to calculate the weighted average of all marginal contributions of each Transformer in an LLM.
-To illustrate the calculation process, we consider a simple LLM with 3 Transformers as shown in Figure 2.
-Since all Transformers have the same structure, we can easily mask a Transformer during inference and do not affect SV calculation, as shown in the bottom right corner of Figure 2.
+Here, we leverage the concept to calculate the weighted average of all marginal contributions of each Transformer layer in an LLM.
+To illustrate the calculation process, we consider a simple LLM with 3 Transformer layers as shown in Figure 3.
+Since all Transformer layers have the same structure, we can easily mask a Transformer layer during inference and do not affect SV calculation, as shown in the bottom right corner of Figure 3.
 We use the accuracy of an LLM as the basis for calculating the SVs.
-Figure 2: Contribution analysis of 3 Transformers in an LLM. For intuitive presentation, we use accuracy to represent the contribution of different Transformers.
-We first identify each Transformer’s contribution when they participate individually, when 2 participate together, and when all 3 participate together.
-Particularly, as an LLM without any Transformer cannot function normally, we assume the accuracy to be 0.
-Then, we consider all possible combinations of Transformers and calculate their marginal values (e.g.
-what value does each Transformer add when Transformer 1 enters the LLM first, followed by Transformer 2, and then Transformer 3).
-Finally, we need to add them up and work out the SV (i.e.
-the average) for each Transformer.
-To formalize this process, an LLM consists of $\mathcal{T}$ Transformers and a value function $\nu$.
-Perplexity (PPL) is usually used to evaluate LLMs in LLM pruning JAISWAL et al.
-(2024b); Xia et al.
-(2024); Ma et al.
-(2023).
+We first identify each Transformer layer’s contribution when they participate individually, when 2 participate together, and when all 3 participate together.
+Particularly, as an LLM without any Transformer layer cannot function normally, we assume the accuracy to be 0.
+Then, we consider all possible combinations of Transformer layers and calculate their marginal values (e.g.
+what value does each Transformer layer add when Transformer layer 1 enters the LLM first, followed by Transformer layer 2, and then Transformer layer 3).
+Finally, we need to add them up and work out the SV (i.e., the average) for each Transformer layer.
+To formalize this process, an LLM consists of $\mathcal{T}$ Transformer layers and a value function $\nu$.
+Perplexity (PPL) is usually used to evaluate LLMs in LLM pruning [31; 32; 9].
 The lower the value, the better the performance.
 Therefore, the value function is denoted as $\nu(\mathcal{T})=1/\text{PPL}(\mathcal{T})$.
-Particularly, if there is no Transformer, the value function must be 0, i.e., $\nu(\emptyset)=0$.
+Particularly, if there is no Transformer layer, the value function must be 0, i.e., $\nu(\emptyset)=0$.
 Let $\mathcal{S}\subseteq\mathcal{T}$ denote a subset of $\mathcal{T}$.
-According to Shapley (1953), the SV of any Transformer $t\in\mathcal{T}$ is:
+According to [33], the SV of any Transformer layer $t\in\mathcal{T}$ is:
 $\phi_{t}=\sum_{\mathcal{S}\subseteq\mathcal{T}\setminus\{t\}}w_{\mathcal{S}}[\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})],\forall t\in\mathcal{T},$ (1)
 where $w_{\mathcal{S}}=T\cdot\binom{T-1}{\mathcal{S}}=\frac{T!}{|\mathcal{S}|!(T-|\mathcal{S}|-1)!}$ is a coefficient.
-$\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})$ is known as Transformer $t$’s marginal contribution.
+$\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})$ is known as Transformer layer $t$’s marginal contribution.
 Eq.
-(1) yields the following desirable properties:
-  • Efficiency: The sum of all Shapley values equals to the total contribution towards an LLM by all Transformers, i.e., $\sum_{t\in\mathcal{T}}\phi_{t}=v(\mathcal{T})$.
-This ensures that all contributions from Transformers are fully allocated with no “leftover” or “missing” gain.
-  • Symmetry: If two Transformers contribute equally to an LLM (i.e., they have the same marginal contribution in all subsets), they should have the same Shapley value, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}\cup\{i\}),\forall\mathcal{S}\subseteq\mathcal{T}\setminus\{t,i\}\Rightarrow\phi_{t}=\phi_{i}$.
-This ensures fairness, as Transformers with the same effect on an LLM are not favored over one another.
-  • Dummy: If a Transformer does not add any marginal contribution to any subset of Transformers, its Shapley value is 0, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}),\forall\mathcal{S}\subseteq\mathcal{T}\Rightarrow\phi_{t}=0$.
-This ensures that Transformers that do not affect an LLM’s output receive no contribution value.
-Eq.
-(1) involves calculating the marginal contribution for every subset $\mathcal{S}$, which results in a combinatorial explosion as the number of Transformers increases.
-Specifically, for $T$ Transformers, the total number of possible subsets $\mathcal{S}$ is $2^{T}$.
-For instance, with 32 Transformers in LLaMa-7B, the number of possible subsets becomes $2962^{32}=4,294,967,$ 296.
-The exact computation of Shapley values requires evaluating every subset that excludes the given Transformer, meaning that $32\times 2^{32}$ evaluations are needed.
+(1) calculaties the marginal contribution for every subset $\mathcal{S}$, which results in a combinatorial explosion as the number of Transformer layers increases.
+Specifically, for $T$ Transformer layers, the total number of possible subsets $\mathcal{S}$ is $2^{T}$.
+For instance, with 32 Transformer layers in LLaMa-7B, the number of possible subsets becomes $2962^{32}=4,294,967,$ 296.
+The exact computation of Shapley values requires evaluating every subset that excludes the given Transformer layer, meaning that $32\times 2^{32}$ evaluations are needed.
 Each additional feature doubles the number of subsets that must be considered, causing an exponential increase in the number of computations.
@@
 4.2 Shapley Value Approximation
-Figure 3: Selection of Transformer coalitions under an LLM with 8 Transformers.
-To address the NP-hard problem, this subsection proposes a sliding window-based Shapley value (SWSV) approximation method to estimate the contribution of Transformers.
-The key is that many coalitions of Transformers are useless and meaningless when evaluating LLMs.
-We consider an LLM with 8 Transformers as shown in Figure 3, where Transformers 1 and 8 are active and Transformers 2-7 are masked.
-We have two limitations: 1) too few active Transformers; and 2) too far apart between active Transformers.
-Specifically, a well-trained LLM relies on maintaining a sufficiently large network structure, even in the presence of sparsity, to preserve its representation capacity and functionality Li et al.
-(2022); Lee et al.
-(2024).
+Figure 4: Selection of Transformer coalitions under an LLM with 8 Transformer layers.
+To address the NP-hard problem, this subsection proposes a SWSV method to estimate the contribution of Transformer layers.
+The key is that many coalitions of Transformer layers are useless and meaningless when evaluating LLMs.
+We consider an LLM with 8 Transformer layers as shown in Figure 4, where Transformers 1 and 8 are active and Transformer layers 2-7 are masked.
+We have two limitations: 1) too few active Transformer layers; and 2) too far apart between active Transformer layers.
+Specifically, a well-trained LLM relies on maintaining a sufficiently large network structure, even in the presence of sparsity, to preserve its representation capacity and functionality [34; 35].
 Sparsity, while effective in reducing computational costs, must be implemented judiciously to ensure that the structural integrity and critical pathways of the model remain intact.
-An LLM with an insufficient number of active Transformer layers—such as configurations involving only two active layers (e.g., Transformers 1 and 8)—is highly likely to suffer from significant performance degradation.
+An LLM with an insufficient number of active Transformer layers—such as only two active layers (e.g., Transformer layers 1 and 8)—is highly likely to suffer from significant performance degradation.
 This occurs because such configurations disrupt the hierarchical processing of information, resulting in incomplete or suboptimal feature representations.
-Moreover, when many intermediate Transformers are masked, the direct transmission of outputs from an earlier Transformer (e.g., Transformer 1) to a much later one (e.g., Transformer 8) bypasses the essential intermediate processing steps.
+Moreover, when many intermediate Transformer layers are masked, the direct transmission of outputs from an earlier Transformer layer (e.g., Transformer layer 1) to a much later one (e.g., Transformer layer 8) bypasses the essential intermediate processing steps.
 This shortcut diminishes the model’s ability to learn and propagate nuanced information through the intermediate layers, leading to a distorted representation of the Shapley value.
 Such scenarios are problematic as they fail to reflect the contributions of individual layers within the model architecture.
-To address these issues, it is critical to design an SV approximation method to avoid extreme patterns with few Transformers.
-The goal is to improve the accuracy of SV estimation, enhance model performance, and reduce computational complexity by preserving critical Transformers.
-Figure 4: Illustration of the proposed approximation.
-To this end, we propose a sliding window-based Shapley value approximation method for evaluating the contribution of Transformers in LLMs (Figure 4).
-When calculating the SV of Transformer $t\in\mathcal{T}$, we consider the sliding window with $N=|\mathcal{S}_{t}|$ Transformers, where these Transformers are closely connected.
-Figure 4 shows the sliding window with 3 Transformers, where the SV of Transformer 1 is only related to the Transformers in this sliding window.
+To address these issues, it is critical to design an SV approximation method to avoid extreme patterns with few Transformer layers.
+The goal is to improve the accuracy of SV estimation, enhance model performance, and reduce computational complexity by preserving critical Transformer layers.
+Figure 5: Illustration of SWSV.
+To this end, we propose a sliding window-based Shapley value approximation method for evaluating the contribution of Transformer layers in LLMs (Figure 5).
+When calculating the SV of Transformer layer $t\in\mathcal{T}$, we consider the sliding window with $N=|\mathcal{S}_{t}|$ Transformer layers, where these Transformer layers are closely connected.
+Figure 5 shows the sliding window with 3 Transformer layers, where the SV of Transformer layer 1 is only related to the Transformer layers in this sliding window.
 It is expressed as:
@@
 where $\mathcal{T}_{t}=\mathcal{T}\setminus\mathcal{S}_{t}$ is coalitions without $\mathcal{S}_{t}$.
-Neighboring Transformers around a target Transformer $t$ play a significant role in determining its SV, as they strongly influence the context and interactions that contribute to the target Transformer’s functionality.
-To effectively capture these interdependencies, a sliding window strategy is employed, which selects a set of neighboring Transformers both before and after the target Transformer.
-This approach ensures that the SV calculation takes into account the most relevant local interactions within the Transformer stack.
-The sliding window size $N$ is generally set to an odd number, such as 3, 5, or 7, to maintain symmetry around the target Transformer.
-For instance, with a sliding window of size 3, the calculation includes one preceding Transformer ($t-1$), the target Transformer itself ($t$), and one succeeding Transformer ($t+1$).
-This balanced selection ensures that contributions from both upstream and downstream Transformers are equally considered, facilitating a more accurate evaluation of the target Transformer’s importance.
-Input: $\mathcal{T}$, $N$, an empty set $\mathcal{V}$, and a dataset;
-Output: Shapley values of Transformers;
-1
-2while $t\in\mathcal{T}$ do
+Input: $\mathcal{T}$, $N$, $\mathcal{V}$, and a dataset;
+Output: Shapley values $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$;
+1 while $t\in\mathcal{T}$ do
 2       Determine the sliding window $\mathcal{S}_{t}$;
@@
 6                   Calculate $\nu(\mathcal{S}\cup\mathcal{T}_{t}\cup\{t\})$;
-8
+7
 8             end if
@@
 11                   Calculate $\nu(\mathcal{S}\cup\mathcal{T}_{t})$;
-13
+12
 13             end if
-15
+14
 15       end for
 16      Calculate $\hat{\phi}_{t}$ by (2);
-18
+17
 18 end while
 Algorithm 1 SWSV
-Special considerations are made for boundary Transformers, such as the first and last ones, where the default sliding window size might result in insufficient neighbors.
-In these cases, the sliding window is adjusted to include the available Transformers, while maintaining the specified number of neighbors as much as possible.
-For example, when evaluating the first Transformer, additional downstream Transformers can be included to compensate for the absence of preceding Transformers.
-Similarly, for the last Transformer, upstream Transformers are prioritized.
-This ensures consistency in the Shapley value calculations across all Transformers, preserving the intended neighborhood size while respecting the structural boundaries of the model.
+Neighboring Transformer layers around a target Transformer layer $t$ play a significant role in determining its SV, as they strongly influence the context and interactions that contribute to the target Transformer layer’s functionality.
+To effectively capture these interdependencies, a sliding window strategy is employed, which selects a set of neighboring Transformer layers both before and after the target Transformer layer.
+This approach ensures that the SV calculation takes into account the most relevant local interactions within the Transformer layer stack.
+The sliding window size $N$ is generally set to an odd number, such as 3, 5, or 7, to maintain symmetry around the target Transformer layer.
+For instance, with a sliding window of size 3, the calculation includes one preceding Transformer layer ($t-1$), the target Transformer layer itself ($t$), and one succeeding Transformer layer ($t+1$).
+This balanced selection ensures that contributions from both upstream and downstream Transformer layers are equally considered, facilitating a more accurate evaluation of the target Transformer layer’s importance.
+Input: an LLM and a dataset;
+Output: The Pruned LLM;
+1 Calculate the contributions $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$ of Transformer layers by SWSV;
+2 Allocate the pruning ratios $\{\rho_{t}\}_{t\in\mathcal{T}}$ by Eq. (3);
+3 while $t\in\mathcal{T}$ do
+4       Prune the Transformer layer $t$ by an advanced pruning method at the pruning ratio $\rho_{t}$;
+5
+6 end while
+7Evaluate the pruned LLM;
+Algorithm 2 SV-NUP
+Special considerations are made for boundary Transformer layers, such as the first and last ones, where the default sliding window size might result in insufficient neighbors.
+In these cases, the sliding window is adjusted to include the available Transformer layers, while maintaining the specified number of neighbors as much as possible.
+For example, when evaluating the first Transformer layer, additional downstream Transformer layers can be included to compensate for the absence of preceding Transformer layers.
+Similarly, for the last Transformer layer, upstream Transformer layers are prioritized.
+This ensures consistency in the Shapley value calculations across all Transformer layers, preserving the intended neighborhood size while respecting the structural boundaries of the model.
 The pseudocode of SWSV is shown in Algorithm 1.
-Its computational complexity is $\mathcal{O}(T2^{N-1})$, which is much lower than the original complexity.
-4.3 SWSV for Pruning LLMs
-After calculating the contributions of Transformers (${\hat{\phi}_{t}},{t\in\mathcal{T}}$), these contributions can be utilized to allocate the pruning ratio for each Transformer.
-Transformers with greater contributions are considered more important and thus have smaller pruning ratios.
+Its computational complexity is $\mathcal{O}(T2^{N-1})$, significantly lower than the original $\mathcal{O}(2^{T})$.
+For instance, computing contributions for LLaMA-7B with $N=5$ requires only 512 subsets, compared to $4,294,967,296$ in the original approach.
+4.3 SV-NUP for LLMs
+After calculating the contributions of Transformer layers (${\hat{\phi}_{t}},{t\in\mathcal{T}}$), these contributions can be utilized to allocate the pruning ratio for each Transformer layer.
+Transformer layers with greater contributions are considered more important and thus have smaller pruning ratios.
 Therefore, the pruning ratio $\rho_{t}$ is inversely proportional to the contribution, i.e., $\rho_{t}\propto 1/\hat{\phi}_{t}$.
-However, since overly sparse Transformers can degrade the performance of LLMs, it is necessary to constrain the allocated pruning ratios.
+However, since overly sparse Transformer layers can degrade the performance of LLMs, it is necessary to constrain the allocated pruning ratios.
 To achieve this, we introduce a small positive value $\lambda$ (e.g., 0.05 or 0.1) as a boundary.
@@
 where $a_{t}=\frac{2\lambda(\hat{\phi}_{t}-\min{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}})}{\max{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}}-\min{\{\hat{\phi}_{t}}\}_{t\in\mathcal{T}}}$.
-SV-NUP is a framework which can be applied to advanced pruning methods (e.g., Wanda Sun et al. (2023) and SparseGPT Frantar and Alistarh (2023)). The pseudocode of SV-NUP is presented in Algorithm 2.
-Input: an LLM and a dataset;
-Output: The Pruned LLM;
-1 Calculate the contributions $\{\hat{\phi}_{t}\}_{t\in\mathcal{T}}$ of Transformers by SWSV;
-2 Allocate the pruning ratios $\{\rho_{t}\}_{t\in\mathcal{T}}$ by Eq. (3);
-3 while $t\in\mathcal{T}$ do
-4       Prune the Transformer $t$ by an advanced pruning method at the pruning ratio $\rho_{t}$;
-5
-6 end while
-7Evaluate the pruned LLM;
-Algorithm 2 SV-NUP
+SV-NUP is a framework which can be applied to advanced pruning methods (e.g., Wanda [14] and SparseGPT [4]). The pseudocode of SV-NUP is presented in Algorithm 2.
 5 Experimental Evalution
-5.1 Experiment Settings
-We evaluate the proposed SV-NUP method on the LLaMA-V1, LLaMA-V2, and OPT model families Frantar and Alistarh (2023).
-Due to computational resource constraints, our focus is on the 7-billion and 13-billion parameter variants.
-All LLMs are quantized with float16 and the sequence length is set as 2048.
-To assess the performance of the pruned LLMs, we employ two general metrics: PPL and zero-shot evaluation Sun et al.
-(2023).
-PPL reflects the model’s ability to predict the next word given the preceding context Muhlgay et al.
-(2023), where lower values indicate better performance.
-We evaluate PPL using the WikiText-2 dataset Merity et al.
-(2016), selecting two randomly downloaded subsets of WikiText-2 for testing.
-For zero-shot evaluation, we choose 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness Mihaylov et al.
-(2018), where higher scores indicate better performance.
-Table 1: PPL↓  performance of LLaMA family pruned by 3 existing methods with/without (w./wo.) SV-NUP at 70% sparsity. We highlight the improved performance in blue.
-LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B
+LLMs and Baselines.
+We evaluate SV-NUP on the LLaMA-V1/v2/v3, and OPT model families [4].
+Due to computational constraints, we select LLMs with up to 13 billion parameters, quantized to float16 with a sequence length of 2048.
+We select 3 baselines for comparison: 1) Magnitude (2017) [36], 2) Wanda (2024) [14], and 3) SparseGPT (2023) [4].
+The proposed SV-NUP method can be integrated into these baselines, allowing us to directly assess its performance gains.
+In addition, we compare SV-NUP with OWL (2024) [12] and ALS (2024) [37], two pruning ratio allocation frameworks based on outliers and information orthogonality, respectively.
+We implement all experiments in PyTorch and use the HuggingFace library to download the pre-trained LLMs and datasets.
+All experiments are conducted on a server equipped with 4 NVIDIA A100 GPUs (40GB each), one AMD EPYC CPU, and 252 GB of memory.
+The pruning ratio constraint $\lambda$ is set to 0.1, and the sliding window size $N$ is chosen as 3, 5, or 7.
+Notably, since ALS does not publicly release its source code, we report its results directly from the original paper.
+Experiment Settings.
+We implement all experiments in PyTorch and use the HuggingFace library to download the pre-trained LLMs and datasets.
+All experiments are performed on a server with 4 NVIDIA A100 GPUs, 1 AMD EPYC CPU, and 252 GB of memory.
+To assess the performance of the pruned LLMs, we employ two general metrics: PPL and zero-shot evaluation [14].
+PPL reflects the model’s ability to predict the next word given the preceding context [38], where lower values indicate better performance.
+We evaluate PPL using the WikiText-2 dataset [39], selecting two randomly downloaded subsets of WikiText-2 for testing.
+For zero-shot evaluation, we choose 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness [40], where higher scores indicate better performance.
+Detailed settings are shown in Appendix B.
+Table 1: PPL↓  performance of LLaMA-v1/v2 family pruned by 3 existing methods with (w.) SV-NUP at 70% sparsity. We highlight the improved performance in blue.
+Model LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA2-13B
 Dense 5.68 5.09 5.47 4.88
 Magnitude 48426.42 84531.15 49808.59 214.21
-SV-NUP w. Magnitude 30586.90 (-36.84%) 10632.55 (-87.42%) 31729.74 (-36.30%) 82.98 (-61.26%)
+Magnitude w. SV-NUP 30586.90 (-36.84%) 10632.55 (-87.42%) 31729.74 (-18.01%) 82.98 (-18.01%)
 Wanda 70.84 41.10 57.20 34.06
-SV-NUP w. Wanda 34.95 (-50.66%) 15.92 (-61.28%) 29.56 (-48.33%) 20.46 (-39.94%)
+Wanda w. SV-NUP 34.95 (-50.66%) 15.92 (-61.28%) 29.56 (-18.01%) 20.46 (-18.01%)
 SparseGPT 18.42 13.74 17.87 14.39
-SV-NUP w. SparseGPT 15.10 (-18.01%) 11.06 (-19.55%) 14.84 (-17.00%) 12.48 (-13.27%)
-5.2 Comparison Baselines
-We select 3 baselines for comparison: 1) Magnitude Jaiswal et al.
-(2024a), 2) Wanda Sun et al.
-(2023), and 3) SparseGPT Frantar and Alistarh (2023).
-The proposed SV-NUP method is integrated into these baselines, allowing us to directly observe the performance improvement brought by it.
-In addition, we compare SV-NUP with OWL Yin et al.
-(2023), a pruning ratio allocation framework based on outliers.
-We implement all experiments in PyTroch and use the HuggingFace library to download the pre-trained LLMs and datasets.
-All experiments are performed on a server with 4 NVIDIA A100 GPUs, 1 AMD EPYC CPU, and 252 GB of memory.
-The limitation $\lambda$ of the pruning ratio is set to 0.1.
-The size of sliding windows $N$ is set as 3, 5, or 7.
-5.3 PPL Improvement by SV-NUP
+SparseGPT w. SV-NUP 15.10 (-18.01%) 11.06 (-19.55%) 14.84 (-18.01%) 12.48 (-18.01%)
+PPL Improvement by SV-NUP.
 We first evaluate the PPL performance gains achieved by integrating SV-NUP into LLM pruning methods, as shown in Table 1.
@@
 Overall, SV-NUP emerges as a crucial enhancement, stabilizing pruning outcomes and delivering superior results, especially when combined with advanced methods like SparseGPT.
-Table 2: PPL↓  performance of LLaMA family and OPT family pruned by different methods at 50% sparsity. We highlight the best performance in bold.
-TinyLLaMA LLaMA-7B Vicuna-7B LLaMA2-7B LLaMA3.2-3B OPT-2.7B OPT-6.7B
-Dense 7.7735 5.6772 6.9031 5.4721 7.8137 12.4705 10.8602
-Magnitude 21.6367 17.2882 24.0034 16.0301 139.4124 265.2033 968.7209
-OWL w. Magnitude 18.7673 16.3453 21.3578 15.7421 99.2163 207.0598 363.6955
-SV-NUP w. Magnitude 18.1645 15.9755 20.7409 15.3144 89.4710 158.2674 308.4234
-Wanda 10.8549 7.0884 8.5325 6.7741 12.6977 13.9611 12.0780
-OWL w. Wanda 11.1533 7.0941 8.5731 6.7954 12.6735 14.7622 12.3921
-SV-NUP w. Wanda 11.0834 7.0610 8.4428 6.7959 12.1578 14.0323 12.0065
-SparseGPT 9.8810 6.8701 8.0767 6.6000 11.0357 12.0591 11.2711
-OWL w. SparseGPT 10.1521 6.8785 8.2247 6.6388 10.9828 13.3490 11.5101
-SV-NUP w. SparseGPT 10.1308 6.8336 8.0762 6.5843 10.8495 12.9768 11.2582
-Table 3: PPL↓  and classification performance of Vicuna-7B and OPT-6.7B pruned by different methods at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
-PPL↓ BoolQ RTE WG OBQA Mean
-Vicuna-7B Dense 6.9031 0.7810 0.6823 0.6938 0.3460 0.6258
-Magnitude 24.0034 0.5462 0.5271 0.5817 0.2200 0.4687
-OWL w. Magnitude 21.3578 0.6067 0.5784 0.6056 0.2300 0.5052
-SV-NUP w. Magnitude 20.7409 0.6037 0.5884 0.5959 0.2360 0.5060
-Wanda 8.5325 0.6969 0.6859 0.6519 0.2960 0.5827
-OWL w. Wanda 8.5731 0.6315 0.7040 0.6685 0.2960 0.5750
-SV-NUP w. Wanda 8.4428 0.6976 0.6968 0.6630 0.2880 0.5863
-SparseGPT 8.0767 0.6437 0.6534 0.6480 0.2940 0.5598
-OWL w. SparseGPT 8.2247 0.6780 0.6715 0.6472 0.2840 0.5702
-SV-NUP w. SparseGPT 8.0762 0.6945 0.6823 0.6567 0.3020 0.5839
-OPT-6.7B Dense 10.8602 0.6606 0.5523 0.6519 0.2760 0.5352
-Magnitude 968.7209 0.3804 0.5271 0.5059 0.1760 0.3974
-OWL w. Magnitude 363.6955 0.3813 0.5271 0.5564 0.2140 0.4197
-SV-NUP w. Magnitude 308.4234 0.3908 0.5271 0.5567 0.2120 0.4217
-Wanda 12.0780 0.6214 0.5271 0.6014 0.2380 0.4970
-OWL w. Wanda 12.3921 0.6220 0.5271 0.5864 0.2380 0.4934
-SV-NUP w. Wanda 12.0065 0.6211 0.5271 0.6069 0.2420 0.4993
-SparseGPT 11.2711 0.6358 0.5343 0.6433 0.2540 0.5168
-OWL w. SparseGPT 11.5101 0.6324 0.5307 0.6393 0.2320 0.5086
-SV-NUP w. SparseGPT 11.2582 0.6410 0.5543 0.6338 0.2540 0.5208
-Table 4: PPL↓  performance under different $N$ at 50% sparsity.
+PPL and Zero-shot Performance.
+The results in Table 2 underscore the effectiveness of integrating SV-NUP, OWL, and ALS into existing pruning methods for the LLaMA and OPT model families at 50% sparsity.
+Particularly, as ALS does not release its source code, the results of ALS are not fair and are for reference only.
+In most scenarios, SV-NUP consistently demonstrates superior performance.
+Specifically, Magnitude combined with SV-NUP consistently outperforms both its standalone version and Magnitude with OWL.
+Notably, while SV-NUP slightly improves the performance of Wanda and SparseGPT, the gains are less pronounced, as the pruned LLMs at 50% sparsity already exhibit robust performance, leaving limited room for further enhancements.
+Interestingly, SV-NUP’s effectiveness is more apparent with larger LLMs, such as the 7-billion-parameter models, whereas its impact diminishes for smaller OPT-2.7B.
+Table 2: PPL↓  performance of LLaMA-v1/v2/v3 family and OPT family pruned by different methods at 50% sparsity. We highlight the best performance in bold. We set the results of ALS as gray.
+Model LLaMA-7B Vicuna-7B LLaMA2-7B LLaMA3.2-3B OPT-2.7B OPT-6.7B
+Dense 5.6772 6.9031 5.4721 7.8137 12.4705 10.8602
+Magnitude 17.2882 24.0034 16.0301 139.4124 265.2033 968.7209
+Magnitude w. ALS 16.8000 15.1900 950.0000
+Magnitude w. OWL 16.3453 21.3578 15.7421 99.2163 207.0598 363.6955
+Magnitude w. SV-NUP 15.9755 20.7409 15.3144 89.4710 158.2674 308.4234
+Wanda 7.0884 8.5325 6.7741 12.6977 13.9611 12.0780
+Wanda w. ALS 12.4700 11.6100 19.1600
+Wanda w. OWL 7.0941 8.5731 6.7954 12.6735 14.7622 12.3921
+Wanda w. SV-NUP 7.0610 8.4428 6.7959 12.1578 14.0323 12.0065
+SparseGPT 6.8701 8.0767 6.6000 11.0357 12.0591 11.2711
+SparseGPT w. ALS 11.8700 10.9900 12.2900
+SparseGPT w. OWL 6.8785 8.2247 6.6388 10.9828 13.3490 11.5101
+SparseGPT w. SV-NUP 6.8336 8.0762 6.5843 10.8495 12.9768 11.2582
+Table 3 compares PPL of pruning methods (Magnitude, Wanda, SparseGPT) combined with ALS, OWL, and SV-NUP across LLaMA-7B, LLaMA2-7B, and OPT-6.7B at 30–50% sparsity.
+SV-NUP achieves the lowest PPL in nearly all cases, demonstrating consistent superiority.
+For example, on OPT-6.7B at 50% sparsity, SV-NUP attains a PPL of 308.4234 (Magnitude) and 12.0065 (Wanda), markedly better than OWL (363.6955, 12.3921) and ALS (950.0000, 19.1600).
+The gains are especially pronounced at higher sparsities and for larger models, highlighting SV-NUP’s robustness and scalability.
+These results solidify SV-NUP as a state-of-the-art pruning technique for efficient model compression.
+Table 3: PPL↓  performance of LLaMA-7B, LLaMA2-7B, and OPT-6.7B at different sparsities.
+Model LLaMA-7B LLaMA2-7B OPT-6.7B
+Sparsity 30% 40% 50% 30% 40% 50% 30% 40% 50%
+Magnitude w. ALS 16.8000 9.6000 11.0300 15.1900 950.0000
+Magnitude w. OWL 6.7413 8.8245 16.3453 6.3403 8.1432 15.7421 12.8504 20.4357 363.6955
+Magnitude w. SV-NUP 6.6769 8.5035 15.9755 6.3438 8.2723 15.3144 12.3753 18.7027 308.4234
+Wanda w. ALS 12.4700 9.1500 9.8100 11.6100 19.1600
+Wanda w. OWL 6.0066 6.3583 7.0941 5.7712 6.0827 6.7954 10.7365 11.2408 12.3921
+Wanda w. SV-NUP 5.9885 6.3294 7.0610 5.7554 6.0702 6.7959 10.6790 11.1051 12.0065
+SparseGPT w. ALS 11.8700 9.1100 9.6700 10.9900 12.2900
+SparseGPT w. OWL 5.9512 6.2543 6.8785 5.7599 6.0340 6.6388 10.9449 11.0805 11.5101
+SparseGPT w. SV-NUP 5.9483 6.2285 6.8336 5.7535 6.0306 6.6243 10.8853 10.9279 11.2582
+Table 4: Zero-shot performance↑  of LLaMA-7B and LLaMA2-7B pruned by different methods at 50% sparsity. The "mean" is the average of BoolQ, RTE, WG, and OBQA.
+Model LLaMA-7B LLaMA2-7B
+Zero-shot task BoolQ RTE WG OBQA Mean↑ BoolQ RTE WG OBQA Mean↑
+Dense 75.08% 66.06% 70.09% 34.20% 61.36% 77.71% 62.82% 69.14% 31.40% 60.27%
+Magnitude 54.56% 54.15% 59.43% 22.60% 47.68% 62.94% 57.04% 63.38% 26.80% 52.54%
+Magnitude w. ALS 59.82% 54.51% 61.25% 36.60% 53.05% 71.38% 55.24% 65.19% 41.40% 58.30%
+Magnitude w. OWL 58.69% 55.23% 61.56% 27.80% 50.82% 63.88% 53.79% 62.90% 28.60% 52.29%
+Magnitude w. SV-NUP 57.63% 57.76% 61.40% 27.20% 51.00% 64.92% 55.96% 64.64% 28.60% 53.53%
+Wanda 72.45% 55.40% 66.61% 28.60% 55.77% 75.81% 54.87% 68.35% 31.20% 57.56%
+Wanda w. ALS 73.70% 60.65% 66.30% 38.60% 59.81% 75.47% 54.87% 67.80% 44.80% 60.74%
+Wanda w. OWL 73.43% 53.43% 65.98% 30.60% 55.86% 77.22% 54.51% 68.51% 30.40% 57.66%
+Wanda w. SV-NUP 72.91% 55.23% 67.56% 29.80% 56.38% 76.18% 54.15% 69.40% 31.20% 57.73%
+SparseGPT 74.43% 49.82% 68.11% 28.20% 55.14% 70.46% 54.51% 67.01% 28.00% 55.00%
+SparseGPT w. ALS 74.28% 54.87% 66.77% 39.00% 58.73% 70.98% 55.96% 67.96% 40.00% 58.73%
+SparseGPT w. OWL 72.84% 54.51% 67.88% 26.00% 55.31% 71.50% 56.32% 68.19% 28.40% 56.10%
+SparseGPT w. SV-NUP 73.30% 54.51% 68.75% 26.80% 55.84% 68.84% 63.54% 67.96% 28.60% 57.23%
+Table 4 presents the zero-shot performance of various pruning methods applied to LLaMA-7B and LLaMA2-7B across four downstream tasks.
+The unpruned dense models serve as baselines, achieving 61.36% and 60.27% average accuracy, respectively.
+Among the pruning strategies, magnitude pruning yields the most significant performance degradation, whereas both Wanda and SparseGPT show stronger resilience, with SparseGPT with SV-NUP achieving up to 57.23% on LLaMA2-7B, closely matching the dense baseline.
+Notably, our proposed method SV-NUP consistently improves performance across all pruning strategies and model backbones.
+It is worth noting that the results of ALS occasionally outperform the dense models; we suspect this is due to differences in experimental settings or hyperparameters rather than true pruning benefits.
+Thus, ALS results are reported for reference only and are excluded from direct comparison.
+Overall, the results validate the effectiveness and robustness of SV-NUP in mitigating performance loss introduced by non-uniform pruning.
+Table 5: PPL↓  under different $N$ at 50% sparsity.
 LLaMA-7B $N=3$ $N=5$ $N=7$
-SV-NUP w. Magnitude 16.0491 16.0081 15.9755
-SV-NUP w. Wanda 7.09322 7.06104 7.09215
-SV-NUP w. SparseGPT 6.8382 6.83358 6.84875
+Magnitude w. SV-NUP 16.0491 16.0081 15.9755
+Wanda w. SV-NUP 7.0932 7.0610 7.0921
+SparseGPT w. SV-NUP 6.8382 6.8336 6.8487
 OPT-6.7B $N=3$ $N=5$ $N=7$
-SV-NUP w. Magnitude 525.407 308.423 661.884
-SV-NUP w. Wanda 12.0469 12.0339 12.0065
-SV-NUP w. SparseGPT 11.2766 11.2661 11.2582
-Table 5: Pruning ratio allocation by two SV approximation methods on LLaMA-7B and OPT-6.7B at 50% sparsity. The ”mean” is the average of BoolQ, RTE, WG, and OBQA.
-PPL↓ BoolQ RTE WG OBQA Mean
+Magnitude w. SV-NUP 525.4074 308.4234 661.8835
+Wanda w. SV-NUP 12.0469 12.0339 12.0065
+SparseGPT w. SV-NUP 11.2766 11.2661 11.2582
+Windows Size Selection.
+Table 5 reports PPL of pruned LLaMA-7B and OPT-6.7B using different pruning strategies with SV-NUP, evaluated under varying window sizes ($N{=}3,5,7$).
+Across both models and all pruning methods, the PPL remains largely stable as the window size changes, indicating that SV-NUP is robust to the choice of window size.
+These results suggest that SV-NUP does not require careful tuning of the window size.
+Table 6: Pruning ratio allocation by two SV approximation methods on LLaMA-7B and OPT-6.7B at 50% sparsity.
+PPL↓ BoolQ RTE WG OBQA Mean↑
 LLaMA-7B SV-NUP w. Magnitude 15.98 0.58 0.58 0.61 0.27 0.5100
@@
 SV w. SparseGPT 11.33 0.63 0.53 0.63 0.25 0.5111
-Figure 5: Distribution of the allocated pruning ratio on LLaMA-7B by 3 methods at 50% sparsity.
-5.4 PPL and Zero-shot Performance by Baselines
-The results in Table 2 underscore the effectiveness of integrating SV-NUP and OWL into existing pruning methods for the LLaMA and OPT model families at 50% sparsity.
-Across most scenarios, SV-NUP consistently demonstrates superior performance.
-Specifically, Magnitude combined with SV-NUP consistently outperforms both its standalone version and Magnitude with OWL.
-Notably, while SV-NUP slightly improves the performance of Wanda and SparseGPT, the gains are less pronounced, as the pruned LLMs at 50% sparsity already exhibit robust performance, leaving limited room for further enhancements.
-Interestingly, SV-NUP’s effectiveness is more apparent with larger LLMs, such as the 7-billion-parameter models, whereas its impact diminishes for smaller models like TinyLLaMA and OPT-2.7B.
-For Wanda, the integration of SV-NUP results in a notable reduction in PPL compared to Wanda alone, highlighting the synergy between SV-NUP and fine-grained pruning strategies.
-OWL also contributes positively, particularly when integrated with SparseGPT and Wanda, though its impact is slightly less consistent than that of SV-NUP.
-Overall, integrating SV-NUP stabilizes and enhances performance across various pruning methods, emerging as a particularly strong addition.
-It consistently ensures competitive results and maintains lower PPL across a wide range of LLM architectures.
-Table 3 presents the PPL and classification performance of OPT-2.7B and OPT-6.7B models pruned by different methods at 50% sparsity.
-For OPT-2.7B, SV-NUP with SparseGPT achieves a mean score of 0.4951, which is slightly lower than SparseGPT alone (0.5039), suggesting that the integration of SV-NUP with SparseGPT may not provide performance gains and could slightly degrade results for smaller OPT models.
-In contrast, for OPT-6.7B, SV-NUP with SparseGPT achieves a mean score of 0.5208, which is comparable to SparseGPT alone (0.5168), indicating that SV-NUP can maintain performance levels when combined with SparseGPT for larger models.
-Table 4 illustrates the impact of varying sliding window sizes ($N$) on performance.
-The results indicate that $N=5$ and $N=7$ achieve lower PPL for both LLaMA-7B and OPT-6.7B.
-However, the choice of $N$ is highly sensitive to different LLMs and pruning methods, highlighting the importance of selecting an appropriate window size tailored to specific scenarios.
-5.5 Different SV Approximation Methods
-We compare our proposed SWSV method with the existing SV approximation method proposed in Kolpaczki et al.
-(2024) to demonstrate its efficiency, as shown in Table 5.
+Different SV Approximation Methods.
+Table 6 shows our proposed SWSV method with the existing SV approximation method in [41] to demonstrate its efficiency.
 For LLaMA-7B, SWSV with SparseGPT achieves the highest mean score (0.5584), slightly surpassing SV with SparseGPT (0.5493), showcasing its ability to enhance performance.
-Similarly, for OPT-6.7B, SWSV with SparseGPT achieves a mean score of 0.5208, outperforming SV with SparseGPT (0.5111).
-These results demonstrate that SWSV not only provides marginal performance improvements over standard SV, particularly when combined with SparseGPT, but also offers lower computational complexity.
-This highlights its potential as a more efficient and effective sparsity allocation approach.
-Figure 5 shows the distribution of allocated pruning ratios for LLaMA-7B at 50% sparsity across three methods: SWSV, SV, and OWL.
-SWSV demonstrates significant variability, particularly in earlier layers, with pruning ratios fluctuating more widely before stabilizing closer to the latter layers.
+These results demonstrate that SWSV not only provides marginal performance improvements over standard SV, but also offers lower computational complexity.
+Figure 6: Activation cosine similarity.
+Activation Cosine Similarity.
+Figure 6 illustrates the activation cosine similarity across Transformer layers between the original LLaMA-7B and its pruned counterparts using Wanda and Wanda w.
+SV-NUP at 70% sparsity.
+Notably, Wanda w.
+SV-NUP consistently yields higher similarity scores than Wanda alone across all layers (0–30), suggesting superior preservation of the original LLaMA-7B activations post-pruning.
+As pruning error increases, cosine similarity correspondingly declines.
+Further insights into SV-NUP’s advantages over other baselines are provided in Figure 7 in Appendix C.
+These findings substantiate the efficacy of SV-NUP in mitigating architectural distortion during pruning.
+Additional experimental results are presented in Appendix C.
 6 Conclusions and Future Work
-In this paper, we propose SV-NUP, a novel approach for non-uniform pruning of LLMs based on the contribution by each Transformer to overall model performance.
-By leveraging the Shapley value, we propose a method to assess the importance of individual Transformers within LLMs.
+In this paper, we propose SV-NUP, a novel approach for non-uniform pruning of LLMs based on the contribution by each Transformer layer to overall model performance.
+By leveraging the Shapley value, we can assess the importance of individual Transformer layers within LLMs.
 To further address the computational complexity associated with the Shapley value, we design a sliding window-based approximation method.
-Extensive experiments have been carried out on LLMs including LLaMA-V1, LLaMA-V2 and OPT model families in comparison with three state-of-the-art pruning methods.
-SV-NUP achieves significant improvements in both PPL and zero-shot performance, demonstrating its promise as a useful LLM pruning method that can better preserve model performance at the given pruning ratios.
+Extensive experiments have been carried out on LLMs, including LLaMA-v1/v2/v3 and OPT model families in comparison with 5 state-of-the-art pruning methods.
+SV-NUP achieves significant improvements in both PPL and zero-shot performance, demonstrating its promise as a useful LLM pruning method that can better preserve model performance.
 In subsequent research, we plan to investigate the integration of pruning and quantization for compressing LLMs to reduce computation and storage.
+Appendix A Extende Related Work
+Uniform Pruning.
+Traditional pruning requires a round of re-training to restore performance, which poses significant challenges for LLMs.
+Researchers have developed pruning algorithms specifically tailored for LLM compression.
+For instance, [9] investigated structured sparse LLMs by applying Taylor pruning to remove entire weight rows, followed by LoRA fine-tuning [13].
+In recent years, the focus has shifted toward unstructured pruning which eliminates the need for fine-tuning.
+SparseGPT [4] employs the Hessian inverse for pruning, followed by weight updates to reduce reconstruction errors between dense and sparse weights.
+Wanda [14] introduced a criterion that incorporates weight magnitude and input activations to preserve outlier features.
+Non-uniform Pruning.
+Uniform layerwise sparsity is commonly used for pruning language models [36, 42], with several studies demonstrating its effectiveness in LLM pruning [43, 44].
+However, there is a growing body of work exploring non-uniform layerwise sparsity, primarily in the context of vision models.
+For example, [15] proposed a non-uniform, scale-free topology inspired by graph theory, which outperforms dense counterparts when applied to restricted Boltzmann machines.
+Subsequent work has improved the scalability of this approach by leveraging Erdős–Rényi graphs [16], extending the method to fully connected layers [11] and convolutional layers [17, 18] to achieve data-free and feedforward-free layerwise sparsity.
+Another approach to non-uniform sparsity involves applying a global threshold across all layers [19, 20, 21, 22].
+However, global pruning has been found to be computationally expensive and ineffective when applied to LLMs.
+Analyzing LLMs.
+The authors in [23] analyzed the contributions of various components in LLMs and their impact on overall performance.
+[24] explored the role of deep layers in LLMs through layer pruning, providing insights into how different layers in relation to model performance.
+[25] examined the redundancy of attention heads in transformer-based models, demonstrating that many attention heads can be pruned without significant performance degradation.
+[26] investigated the behavior of individual attention heads in BERT, revealing that each head serves a distinct role in capturing different linguistic features.
+Probing techniques are widely used to analyze the internal representations of LLMs.
+For instance, [27] employed probing tasks to examine the linguistic information captured by BERT, finding that different layers encode distinct types of linguistic features.
+Furthermore, [28] introduced a suite of probes to analyze the representations learned by contextualized word embeddings, offering insights into how syntactic and semantic information is distributed across layers.
+Recently, [29] used Shapley values to evaluate the importance of layers in LLMs, providing a faithful assessment of their contributions.
+However, these studies have not explored SV-based non-uniform pruning of LLMs.
+SV-NUP bridges this gap.
+Appendix B Implementation Details
+We follow existing works [45, 4, 46] and prune all linear layers in the FFN and MHA modules of LLMs.
+For calibration data, we use WikiText-2-v1, specifically "train-00000-of-00001.parquet" and "test-00000-of-00001.parquet" (https://huggingface.co/datasets/Salesforce/wikitext/tree/main/wikitext-2-v1), randomly selecting 32 segments of 2028 tokens from the trainset to ensure zero-shot pruning with generic internet text.
+For zero-shot evaluation, we adopt 4 tasks—BoolQ, RTE, WinoGrande (WG), and OpenBookQA (OBQA)—from the EleutherAI LM Harness framework modified by [45], enabling robust evaluation of pruned LLMs.
+For SV-NUP, we use 32 segments of 2028 tokens from the WikiText-2-v1 trainset to compute Transformer layer contributions, with the sliding window size $N\in\{3,5,7\}$
+For the SV approximation method [41], the budget and step size are set to $5\times 10^{3}$ and 100, respectively.
+For OWL [12], the outlier parameter is 5, and both OWL and SV-NUP share the same $\lambda$.
+For ALS [37], since the original implementation has not been open-sourced, we directly report the results from the original paper as a reference.
+However, due to potential differences in experimental settings and the lack of reproducibility, these results are not strictly comparable to ours and should be considered as indicative rather than fully fair baselines.
+Appendix C More Experimental Results
+(a) LLaMA-7B with Mean
+(b) LLaMA-7B with Std
+(c) Transformer Layer
+Figure 7: Magnitude distribution of LLaMA-7B and pruning rate distribution allocated by the 3 methods at 50% sparsity.
+Analysis of Layer-wise Pruning Ratios.
+Figure 7 visualizes the allocated pruning ratios across Transformer layers for LLaMA-7B (50% sparsity) under SWSV, SV, and OWL.
+SWSV exhibits greater variability in early layers (e.g., Layer 5–15), with pruning ratios fluctuating ±15% around the target sparsity before stabilizing in later layers.
+This contrasts with OWL’s near-uniform distribution and SV’s intermediate behavior.
+From the magnitude distribution in Figures 7(a) and 7(b), SWSV’s dynamic allocation better aligns with layer sensitivity trends—aggressively pruning robust early layers while preserving critical later ones.
+Table 7: PPL↓  and Zero-shot↑  performance of LLaMA-v1 pruned by different methods at 50% sparsity.
+PPL↓ BoolQ RTE WG OBQA Mean↑
+LLaMA-7B Dense 5.6772 75.08% 66.06% 70.09% 34.20% 61.36%
+Magnitude 17.2882 54.56% 54.15% 59.43% 22.60% 47.68%
+Magnitude w. OWL 16.3453 58.69% 55.23% 61.56% 27.80% 50.82%
+Magnitude w. SV-NUP 15.9755 57.63% 57.76% 61.40% 27.20% 51.00%
+Wanda 7.0884 72.45% 55.40% 66.61% 28.60% 55.77%
+Wanda w. OWL 7.0941 73.43% 53.43% 65.98% 30.60% 55.86%
+Wanda w. SV-NUP 7.0610 72.91% 55.23% 67.56% 29.80% 56.38%
+SparseGPT 6.8701 74.43% 49.82% 68.11% 28.20% 55.14%
+SparseGPT w. OWL 6.8785 72.84% 54.51% 67.88% 26.00% 55.31%
+SparseGPT w. SV-NUP 6.8336 73.30% 54.51% 68.75% 26.80% 55.84%
+Vicuna-7B Dense 6.9031 78.10% 68.23% 69.38% 34.60% 62.58%
+Magnitude 24.0034 54.62% 52.71% 58.17% 22.00% 46.87%
+Magnitude w. OWL 21.3578 60.67% 57.84% 60.56% 23.00% 50.52%
+Magnitude w. SV-NUP 20.7409 60.37% 58.84% 59.59% 23.60% 50.60%
+Wanda 8.5325 69.69% 68.59% 65.19% 29.60% 58.27%
+Wanda w. OWL 8.5731 63.15% 70.40% 66.85% 29.60% 57.50%
+Wanda w. SV-NUP 8.4428 69.76% 69.68% 66.30% 28.80% 58.63%
+SparseGPT 8.0767 64.37% 65.34% 64.80% 29.40% 55.98%
+SparseGPT w. OWL 8.2247 67.80% 67.15% 64.72% 28.40% 57.02%
+SparseGPT w. SV-NUP 8.0762 69.45% 68.23% 65.67% 30.20% 58.39%
+Table 8: PPL↓  and Zero-shot↑  performance of LLaMA-v2/v3 pruned by different methods at 50% sparsity.
+PPL↓ BoolQ RTE WG OBQA Mean↑
+LLaMA2-7B Dense 5.4721 77.71% 62.82% 69.14% 31.40% 60.27%
+Magnitude 16.0301 62.94% 57.04% 63.38% 26.80% 52.54%
+Magnitude w. OWL 15.7421 63.88% 53.79% 62.90% 28.60% 52.29%
+Magnitude w. SV-NUP 15.3144 64.92% 55.96% 64.64% 28.60% 53.53%
+Wanda 6.7741 75.81% 54.87% 68.35% 31.20% 57.56%
+Wanda w. OWL 6.7954 77.22% 54.51% 68.51% 30.40% 57.66%
+Wanda w. SV-NUP 6.7959 76.18% 54.15% 69.40% 31.20% 57.73%
+SparseGPT 6.6000 70.46% 54.51% 67.01% 28.00% 55.00%
+SparseGPT w. OWL 6.6388 71.50% 56.32% 68.19% 28.40% 56.10%
+SparseGPT w. SV-NUP 6.5843 68.84% 63.54% 67.96% 28.60% 57.23%
+LLaMA3.2-3B Dense 7.8137 73.27% 54.51% 69.77% 31.20% 57.19%
+Magnitude 139.4124 42.02% 53.43% 53.43% 14.20% 40.77%
+Magnitude w. OWL 99.2163 41.28% 51.62% 54.78% 16.80% 41.12%
+Magnitude w. SV-NUP 89.4710 44.07% 51.99% 56.99% 17.60% 42.66%
+Wanda 12.6977 66.02% 55.23% 62.51% 25.60% 52.34%
+Wanda w. OWL 12.6735 63.98% 53.07% 64.56% 24.80% 51.60%
+Wanda w. SV-NUP 12.1578 69.88% 55.96% 65.27% 24.40% 53.88%
+SparseGPT 11.0357 70.49% 52.71% 64.88% 24.40% 53.12%
+SparseGPT w. OWL 10.9828 70.28% 49.46% 66.30% 24.80% 52.71%
+SparseGPT w. SV-NUP 10.8495 70.46% 51.62% 65.98% 25.20% 53.32%
+Table 9: PPL↓  and Zero-shot↑  performance of OPT pruned by different methods at 50% sparsity.
+PPL↓ BoolQ RTE WG OBQA Mean↑
+OPT-2.7B Dense 12.4705 60.37% 54.87% 60.77% 25.00% 50.25%
+Magnitude 265.2033 39.66% 52.35% 53.43% 20.40% 41.46%
+Magnitude w. OWL 207.0598 38.17% 52.71% 53.67% 21.20% 41.44%
+Magnitude w. SV-NUP 158.2674 38.59% 52.35% 53.51% 22.40% 41.71%
+Wanda 13.9611 62.26% 51.99% 57.70% 20.40% 48.09%
+Wanda w. OWL 14.7622 61.90% 52.35% 57.70% 20.80% 48.18%
+Wanda w. SV-NUP 14.0042 62.29% 51.62% 58.41% 21.60% 48.48%
+SparseGPT 12.0591 60.70% 54.51% 61.33% 25.00% 50.39%
+SparseGPT w. OWL 13.3490 62.87% 51.62% 58.80% 23.20% 49.12%
+SparseGPT w. SV-NUP 12.9768 63.30% 52.71% 59.43% 22.60% 49.51%
+OPT-6.7B Dense 10.8602 66.06% 55.23% 65.19% 27.60% 53.52%
+Magnitude 968.7209 38.04% 52.71% 50.59% 17.60% 39.74%
+Magnitude w. OWL 363.6955 38.13% 52.71% 55.64% 21.40% 41.97%
+Magnitude w. SV-NUP 308.4234 39.08% 52.71% 55.67% 21.20% 42.17%
+Wanda 12.0780 62.14% 52.71% 60.14% 23.80% 49.70%
+Wanda w. OWL 12.3921 62.20% 52.71% 58.64% 23.80% 49.34%
+Wanda w. SV-NUP 12.0065 62.11% 52.71% 60.69% 24.20% 49.93%
+SparseGPT 11.2711 63.58% 53.43% 64.33% 25.40% 51.68%
+SparseGPT w. OWL 11.5101 63.24% 53.07% 63.93% 23.20% 50.86%
+SparseGPT w. SV-NUP 11.2582 64.10% 55.43% 63.38% 25.40% 52.08%
+Table 10: PPL↓  and Zero-shot↑  performance of LLaMA-7B at different sparsities.
+LLaMA-7B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 6.7413 71.68% 57.04% 69.53% 31.60% 57.46%
+Magnitude w. SV-NUP 6.6769 71.96% 61.73% 68.98% 30.60% 58.32%
+Wanda w. OWL 6.0066 76.21% 62.45% 68.90% 32.80% 60.09%
+Wanda w. SV-NUP 5.9885 75.81% 63.18% 69.30% 32.80% 60.27%
+SparseGPT w. OWL 5.9512 75.57% 62.45% 69.77% 30.80% 59.65%
+SparseGPT w. SV-NUP 5.9483 75.72% 63.18% 70.01% 32.60% 60.38%
+40% Magnitude w. OWL 8.8245 67.03% 56.68% 66.69% 31.20% 55.40%
+Magnitude w. SV-NUP 8.5035 67.34% 59.21% 67.32% 31.00% 56.22%
+Wanda w. OWL 6.3583 74.50% 59.21% 69.30% 31.00% 58.50%
+Wanda w. SV-NUP 6.3294 74.37% 60.65% 68.67% 30.40% 58.52%
+SparseGPT w. OWL 6.2543 74.74% 61.37% 68.82% 30.20% 58.78%
+SparseGPT w. SV-NUP 6.2285 74.71% 59.93% 69.61% 29.80% 58.51%
+50% Magnitude w. OWL 16.3453 58.69% 55.23% 61.56% 27.80% 50.82%
+Magnitude w. SV-NUP 15.9755 57.63% 57.76% 61.40% 27.20% 51.00%
+Wanda w. OWL 7.0941 73.43% 53.43% 65.98% 30.60% 55.86%
+Wanda w. SV-NUP 7.0610 72.91% 55.23% 67.56% 29.80% 56.38%
+SparseGPT w. OWL 6.8785 72.84% 54.51% 67.88% 26.00% 55.31%
+SparseGPT w. SV-NUP 6.8336 73.30% 54.51% 68.75% 26.80% 55.84%
+Table 11: PPL↓  and Zero-shot↑  performance of Vicuna-7B at different sparsities.
+Vicuna-7B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 8.5803 75.41% 67.15% 66.85% 32.00% 60.35%
+Magnitude w. SV-NUP 8.1363 75.69% 69.68% 66.14% 32.20% 60.93%
+Wanda w. OWL 7.3325 74.40% 64.62% 68.75% 33.00% 60.19%
+Wanda w. SV-NUP 7.3159 74.07% 65.70% 68.51% 33.40% 60.42%
+SparseGPT w. OWL 7.2667 74.13% 63.90% 68.67% 31.40% 59.52%
+SparseGPT w. SV-NUP 7.2622 74.31% 64.98% 68.03% 31.80% 59.78%
+40% Magnitude w. OWL 11.8580 69.60% 62.45% 64.48% 30.20% 56.69%
+Magnitude w. SV-NUP 10.8052 73.76% 63.18% 64.48% 29.40% 57.71%
+Wanda w. OWL 7.7982 70.52% 63.90% 66.77% 32.00% 58.30%
+Wanda w. SV-NUP 7.6862 73.15% 67.87% 67.25% 32.60% 60.22%
+SparseGPT w. OWL 7.5926 63.21% 68.23% 67.56% 32.20% 57.80%
+SparseGPT w. SV-NUP 7.5080 71.10% 63.54% 66.69% 31.00% 58.08%
+50% Magnitude w. OWL 21.3578 60.67% 57.84% 60.56% 23.00% 50.52%
+Magnitude w. SV-NUP 20.7409 60.37% 58.84% 59.59% 23.60% 50.60%
+Wanda w. OWL 8.5731 63.15% 70.40% 66.85% 29.60% 57.50%
+Wanda w. SV-NUP 8.4428 69.76% 69.68% 66.30% 28.80% 58.63%
+SparseGPT w. OWL 8.2247 67.80% 67.15% 64.72% 28.40% 57.02%
+SparseGPT w. SV-NUP 8.0762 69.45% 68.23% 65.67% 30.20% 58.39%
+Table 12: PPL↓  and Zero-shot↑  performance of LLaMA2-7B at different sparsities.
+LLaMA2-7B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 6.3403 73.33% 56.32% 69.93% 31.20% 57.69%
+Magnitude w. SV-NUP 6.3438 72.87% 58.12% 70.40% 32.00% 58.35%
+Wanda w. OWL 5.7712 76.79% 57.40% 68.82% 33.20% 59.05%
+Wanda w. SV-NUP 5.7554 77.09% 57.40% 69.38% 33.40% 59.32%
+SparseGPT w. OWL 5.7599 77.55% 57.40% 69.85% 33.80% 59.65%
+SparseGPT w. SV-NUP 5.7535 77.25% 61.01% 69.53% 33.00% 60.20%
+40% Magnitude w. OWL 8.2723 69.94% 57.40% 68.90% 31.20% 56.86%
+Magnitude w. SV-NUP 8.1432 69.91% 57.76% 68.67% 32.20% 57.13%
+Wanda w. OWL 6.0827 75.81% 54.15% 68.75% 32.20% 57.73%
+Wanda w. SV-NUP 6.0702 75.75% 59.57% 69.06% 32.00% 59.09%
+SparseGPT w. OWL 6.0340 75.93% 57.04% 68.51% 31.80% 58.32%
+SparseGPT w. SV-NUP 6.0306 75.54% 56.68% 68.43% 30.80% 57.86%
+50% Magnitude w. OWL 15.7421 63.88% 53.79% 62.90% 28.60% 52.29%
+Magnitude w. SV-NUP 15.3144 64.92% 55.96% 64.64% 28.60% 53.53%
+Wanda w. OWL 6.7954 77.22% 54.51% 68.51% 30.40% 57.66%
+Wanda w. SV-NUP 6.7959 76.18% 54.15% 69.40% 31.20% 57.73%
+SparseGPT w. OWL 6.6388 71.50% 56.32% 68.19% 28.40% 56.10%
+SparseGPT w. SV-NUP 6.5843 68.84% 63.54% 67.96% 28.60% 57.23%
+Table 13: PPL↓  and Zero-shot↑  performance of LLaMA3.2-3B at different sparsities.
+LLaMA3.2-3B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 9.9054 64.71% 57.40% 66.93% 30.20% 54.81%
+Magnitude w. SV-NUP 9.9530 68.20% 57.40% 68.11% 30.00% 55.93%
+Wanda w. OWL 8.4789 71.87% 50.54% 70.01% 31.60% 56.00%
+Wanda w. SV-NUP 8.4111 73.30% 52.35% 70.40% 31.40% 56.86%
+SparseGPT w. OWL 8.3514 73.76% 54.15% 70.09% 31.80% 57.45%
+SparseGPT w. SV-NUP 8.3202 73.67% 54.51% 70.24% 31.00% 57.36%
+40% Magnitude w. OWL 15.8035 54.28% 54.87% 63.38% 25.40% 49.48%
+Magnitude w. SV-NUP 16.2707 53.85% 53.79% 64.25% 24.80% 49.17%
+Wanda w. OWL 9.6285 68.44% 55.96% 67.72% 29.00% 55.28%
+Wanda w. SV-NUP 9.4400 70.24% 55.23% 68.11% 30.00% 55.90%
+SparseGPT w. OWL 9.1586 74.01% 49.46% 69.53% 29.20% 55.55%
+SparseGPT w. SV-NUP 9.0932 71.99% 51.99% 68.43% 30.00% 55.60%
+50% Magnitude w. OWL 99.2163 41.28% 51.62% 54.78% 16.80% 41.12%
+Magnitude w. SV-NUP 89.4710 44.07% 51.99% 56.99% 17.60% 42.66%
+Wanda w. OWL 12.6735 63.98% 53.07% 64.56% 24.80% 51.60%
+Wanda w. SV-NUP 12.1578 69.88% 55.96% 65.27% 24.40% 53.88%
+SparseGPT w. OWL 10.9828 70.28% 49.46% 66.30% 24.80% 52.71%
+SparseGPT w. SV-NUP 10.8495 70.46% 51.62% 65.98% 25.20% 53.32%
+Table 14: PPL↓  and Zero-shot↑  performance of OPT-2.7B at different sparsities.
+OPT-2.7B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 15.0581 46.02% 54.51% 57.70% 24.00% 45.56%
+Magnitude w. SV-NUP 14.4552 48.87% 52.71% 59.67% 22.60% 45.96%
+Wanda w. OWL 12.3628 65.50% 51.26% 60.46% 24.20% 50.36%
+Wanda w. SV-NUP 12.2264 66.27% 51.26% 60.06% 24.20% 50.45%
+SparseGPT w. OWL 12.2349 64.28% 52.35% 60.30% 24.60% 50.38%
+SparseGPT w. SV-NUP 12.1657 66.54% 53.79% 60.14% 22.80% 50.82%
+40% Magnitude w. OWL 22.9237 44.65% 52.35% 58.17% 22.40% 44.39%
+Magnitude w. SV-NUP 22.8636 44.01% 51.99% 58.33% 23.80% 44.53%
+Wanda w. OWL 13.0116 62.48% 52.71% 58.72% 23.60% 49.38%
+Wanda w. SV-NUP 12.7347 62.97% 51.62% 58.80% 24.00% 49.35%
+SparseGPT w. OWL 12.5431 64.62% 53.07% 59.19% 23.40% 50.07%
+SparseGPT w. SV-NUP 12.3850 65.93% 53.07% 59.67% 24.00% 50.67%
+50% Magnitude w. OWL 207.0598 38.17% 52.71% 53.67% 21.20% 41.44%
+Magnitude w. SV-NUP 158.2674 38.59% 52.35% 53.51% 22.40% 41.71%
+Wanda w. OWL 14.7622 61.90% 52.35% 57.70% 20.80% 48.18%
+Wanda w. SV-NUP 14.0042 62.29% 51.62% 58.41% 21.60% 48.48%
+SparseGPT w. OWL 13.3490 62.87% 51.62% 58.80% 23.20% 49.12%
+SparseGPT w. SV-NUP 12.9768 63.30% 52.71% 59.43% 22.60% 49.51%
+Table 15: PPL↓  and Zero-shot↑  performance of OPT-6.7B at different sparsities.
+OPT-6.7B PPL↓ BoolQ RTE WG OBQA Mean↑
+30% Magnitude w. OWL 12.8504 49.79% 53.43% 59.98% 26.00% 47.30%
+Magnitude w. SV-NUP 12.3753 52.78% 53.79% 60.06% 26.20% 48.21%
+Wanda w. OWL 10.7365 65.72% 54.15% 63.93% 27.60% 52.85%
+Wanda w. SV-NUP 10.6790 66.48% 53.43% 64.01% 28.60% 53.13%
+SparseGPT w. OWL 10.9449 67.77% 53.79% 63.77% 27.00% 53.08%
+SparseGPT w. SV-NUP 10.8853 67.89% 54.15% 63.77% 27.00% 53.20%
+40% Magnitude w. OWL 20.4357 43.00% 52.71% 58.48% 25.40% 44.90%
+Magnitude w. SV-NUP 18.7027 44.34% 52.71% 58.72% 25.60% 45.34%
+Wanda w. OWL 11.2408 62.35% 52.35% 62.90% 26.80% 51.10%
+Wanda w. SV-NUP 11.1051 62.94% 52.35% 62.67% 26.20% 51.04%
+SparseGPT w. OWL 11.0805 64.95% 53.79% 64.25% 26.60% 52.40%
+SparseGPT w. SV-NUP 10.9279 66.36% 54.51% 63.85% 26.00% 52.68%
+50% Magnitude w. OWL 363.6955 38.13% 52.71% 55.64% 21.40% 41.97%
+Magnitude w. SV-NUP 308.4234 39.08% 52.71% 55.67% 21.20% 42.17%
+Wanda w. OWL 12.3921 62.20% 52.71% 58.64% 23.80% 49.34%
+Wanda w. SV-NUP 12.0065 62.11% 52.71% 60.69% 24.20% 49.93%
+SparseGPT w. OWL 11.5101 63.24% 53.07% 63.93% 23.20% 50.86%
+SparseGPT w. SV-NUP 11.2582 64.10% 55.43% 63.38% 25.40% 52.08%